### v1

# QUITE: A Query Rewrite System Beyond Rules with LLM Agents

Abstract.
Query rewrite transforms SQL queries into semantically equivalent forms that run more efficiently. Existing approaches mainly rely on predefined rewrite rules, but they handle a limited subset of queries and can cause performance regressions. This limitation stems from three challenges of rule-based query rewrite: (1) it is hard to discover and verify new rules, (2) fixed rewrite rules do not generalize to new query patterns, and (3) some rewrite techniques cannot be expressed as fixed rules. Motivated by the fact that human experts exhibit significantly better rewrite ability but suffer from scalability, and Large Language Models (LLMs) have demonstrated nearly human-level semantic and reasoning abilities, we propose a new approach of using LLMs to rewrite SQL queries beyond rules. Due to the hallucination problems in LLMs, directly applying LLMs often leads to nonequivalent and suboptimal queries. To address this issue, we propose QUITE (query rewrite), a training-free and feedback-aware system based on LLM agents that rewrites SQL queries into semantically equivalent forms with significantly better performance, covering a broader range of query patterns and rewrite strategies compared to rule-based methods. Firstly, we design a multi-agent framework controlled by a finite state machine (FSM) to equip LLMs with the ability to use external tools and enhance the rewrite process with real-time database feedback. Secondly, we develop a rewrite middleware to enhance the ability of LLMs to generate optimized query equivalents. Finally, we employ a novel hint injection technique to improve execution plans for rewritten queries. Extensive experiments show that QUITE reduces query execution time by up to 35.8 % over state-of-the-art approaches and produces 24.1 % more rewrites than prior methods, covering query cases that earlier systems did not handle.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
1. Introduction
Inefficient SQL queries remain a long-standing challenge in many database applications. These inefficiencies often arise from poorly written queries, which are typically produced by inexperienced users or automatically generated by systems (Li, 2019; Zhou et al., 2020; Pirahesh et al., 1992). Query rewrite is a promising approach that transforms an original SQL query into a semantically equivalent form with significantly better performance (Finance and Gardarin, 1991a; Pirahesh et al., 1992; Wang et al., 2022). An effective rewrite can improve the execution time of a poorly written SQL query by several orders of magnitude (Wang et al., 2022; Zhou et al., 2021).
Existing query rewrite approaches typically rely on predefined rewrite rules (Begoli et al., 2018; Graefe and McKenna, 1993; PostgreSQL Global Development Group, 2024; Zhou et al., 2021, 2023; Bai et al., 2023; Li et al., 2024; Sun et al., 2024). Heuristic-based methods apply these rules in a fixed order derived from practical experience (Begoli et al., 2018; Graefe and McKenna, 1993; PostgreSQL Global Development Group, 2024). However, such fixed orders often fail to generalize across different environments and query patterns. Thus, machine learning (ML)-based methods train models on rewrite histories to predict which rules to use and the order in which to apply them (Zhou et al., 2021, 2023). Recently, some approaches have explored the use of Large Language Models (LLMs) to select rewrite rules and determine their application order based on domain knowledge (Li et al., 2024; Sun et al., 2024).
Figure 1. An Example of Using Strategies beyond Rewrite Rules and Using Context-Aware Analysis. (b) boosts speed 37.5x over (a) by using a common table expression (CTE) to eliminate redundant calculations. (c) leverages data characteristics by joining sale (10 M rows) with prod (100 K rows) yields a 100 K-row intermediate (vs. 1.1 M rows in (b) ), achieving a 2.79× speedup.
Figure 2. An Example of Using Query Intent
However, even state-of-the-art query rewrite approaches can handle only a small subset of queries and may lead to performance regressions (Zhou et al., 2021; Li et al., 2024; Sun et al., 2024). This limitation stems from the intrinsic challenges of rule-based query rewrite. First, the rich features of SQL and the nuances of semantics make it rather difficult to discover and verify new rewrite rules (Pirahesh et al., 1992; Chu et al., 2017b; Wang et al., 2022). Even state-of-the-art rule discovery system (Wang et al., 2022) can only support a narrow subset of operators. Second, fixed rewrite rules rely on pattern matching and therefore are fundamentally unable to optimize unseen or complex query patterns (Leis et al., 2015; Bai et al., 2023; Dong et al., 2023). Third, many effective rewrite strategies cannot be expressed by rewrite rules (Dong et al., 2023). For example, the Common Table Expressions (CTE) conversion in Example 1.1 cannot be captured by rewrite rules. This is because such conversion requires explicit control over evaluation order and sub-plan sharing, which is beyond the scope of relational-algebra rewrite rules (Graefe, 1993; Wang et al., 2022; Begoli et al., 2018).
Given these limitations, we revisit the query rewrite problem from the perspective of human experts and identify three key capabilities that humans exhibit, but rule-based systems lack. (1) Humans can perform rewrites that follow rules but go beyond the expressiveness of pattern matching (Example 1.1). (2) Experts can apply intent-oriented rewrites that capture the underlying goal of the query, which are not easily expressed as explicit rules (Example 1.2). (3) Humans can conduct context-aware rewrites by considering the underlying data distributions and cost estimates provided by the query optimizer (Example 1.3).
Example 0.
Query Rewrite beyond Rewrite Rules. Figure 1 (a) is a simplified query from the DSB benchmark (Ding et al., 2021). To avoid repeated calculation of nested subqueries, Figure 1 (b) rewrites them using Common Table Expressions (CTEs). This enables the database engine to compute intermediate results once and reuse them, leading to a 37.5 $\times$ reduction in query latency in our experimental setup.
Example 0.
Query Rewrite based on Query Intent. Figure 2 (a) shows a real-world query from Apache Calcite (apa, 2021). Rule-based rewrite cannot optimize this query (Figure 2 (b). Identifying that the query intent is to sum salaries directly from the EMP table, human can simplify the original query by eliminating unnecessary joins and sub-queries that do not contribute to the intent of the query, achieving a 2.17 $\times$ execution improvement in our experimental setup.
Example 0.
Context-Aware Query Rewrite. In figure 1 (c), instead of directly joining two large tables (the sale with 10 M rows and the ttime with 1.4 M rows), the query first joins the sale table with the prod table (100 K rows). This highly selective join would produce an intermediate result of only  100 K rows. Next we join this with ttime table and this yields a 2.79 $\times$ speedup over (b) in our experimental setup. Such rewrites, which reduce intermediate result size, are rare in rule-based systems and require analyzing data distribution and selectivity.
Although human experts can produce significantly better SQL rewrites than rule-based methods, manual rewrite suffers from poor scalability, especially when there are thousands or even more poorly written queries to optimize in cloud environments. Recently, LLMs have demonstrated nearly human-level semantic and reasoning abilities, making them a natural choice for query rewrite. However, this task remains challenging due to the hallucination problem of LLMs: they may generate responses that are plausible but factually incorrect or semantically invalid. Specifically, there are two fundamental challenges. C1. Ensuring Equivalent Rewrites. Producing equivalent rewrites requires a precise understanding of the rich features of SQL and the nuances of semantics. Even the most powerful LLMs produce rewrites with syntax or semantic errors in nearly 20% of cases (Section 7). C2. Ensuring Optimized Rewrites. Determining whether a rewritten query improves performance depends on multiple factors, including query intent, data distribution, optimizer estimates, system states, index configurations, and others. These contextual factors are typically unavailable to LLMs, and enabling LLMs to effectively utilize such information remains a non-trivial problem (Lao et al., 2024a; Bai et al., 2024; Tan et al., 2025).
To address the limitations of rule-based query rewrite, the poor scalability of manual rewrite, and the unreliable nature of LLMs, we propose QUITE, a training-free and feedback-aware system that smartly leverages LLMs to rewrite SQL queries into semantically equivalent forms with significantly improved performance. QUITE  supports a broader range of query patterns and rewrite strategies, while maintaining good scalability. First, to enable LLMs to capture query data characteristics and leverage database feedback, we design a multi-agent framework controlled by a finite state machine (FSM) to iteratively refine the rewrite process. We decompose the complex rewrite process into a set of subtasks, each handled by a specialized LLM agent to reduce hallucinations and improve rewrite ability. Second, to enhance the ability of LLMs to generate high-quality query rewrites, we provide a rewrite middleware that includes: (1) a structured knowledge base extracted from official documentation and query rewrite discussions on web forums, (2) a hybrid SQL corrector to ensure query equivalence, and (3) an agent memory buffer that manages important rewrite contexts. Third, we employ a novel hint injection technique to generate better execution plans for rewritten queries. This ensures that query execution exactly follows the optimized plan recommended by LLM agents, avoiding unintended modifications by the query optimizer due to inaccurate cost estimations. Additionally, this technique enables further improvements by incorporating physical-level optimizations.
We extensively evaluate QUITE against state-of-the-art query rewrite systems on widely used benchmarks (e.g., TPC-H, DSB, and Calcite) and show that our system delivers up to a 21.9 % speedup in query execution performance over the best-performing alternative. Moreover, it delivers 24.1 % additional rewrites compared to that alternative, covering query cases that previous systems could not handle. In summary, we make the following contributions:
  • We propose QUITE, a training-free and feedback-aware query rewrite system that leverages LLM agents to support a broader range of query patterns and rewrite strategies than existing methods (Section 3).
  • We develop a rewrite middleware including a structured knowledge base, a hybrid SQL corrector, and an agent memory buffer to enhance the rewrite capability of LLM agents (Section 5).
  • We design a multi-agent framework controlled by a finite state machine to enable LLMs to use external tools and enhance the rewrite process with database feedback (Section 4).
  • We propose a hint injection technique to produce better execution plans for rewritten queries. (Section 6)
  • We conduct extensive experiments to demonstrate that our system significantly surpasses state-of-the-art methods in both query performance and query coverage (Section 7).
2. Preliminaries and Related Work
Figure 3. System Overview of QUITE
2.1. Query Rewrite
We first define query rewrite in the realm of database systems:
Definition 2.1 (Query Rewrite).
Query rewrite is a process of transforming a query into another query that is semantically equivalent but with improved performance. It is a distinct step prior to query optimization, operating at the level of the user-accessible application programming interface (API) (e.g., SQL).
Definition 2.2 (Query Equivalence).
Two queries are considered equivalent if they produce the same result for any valid instance of the database schema.
Conventional studies treat query rewriting as a transformation between plain SQL statements, typically using rule-based methods. There are two research directions:
(1) Discovery of new rules. Approaches such as WeTune (Wang et al., 2022) automatically generate and verify logical plan transformations, but they are limited to certain operator types or algebraic equivalences. Techniques based on equality saturation (Nandi et al., 2021) and user-guided tools like QueryBooster (Bai et al., 2023) lack full autonomy and broad applicability. More recently, GENRewrite (Liu and Mozafari, 2024) uses LLMs to suggest rules in natural language. Yet GENRewrite must overcome challenges in formally verifying equivalence, preventing incorrect suggestions, and choosing the best candidates.
(2) Effective use of existing rules. A rule-based rewrite framework must answer two key questions: which rules to include in the rule set and how to select among applicable rules during rewriting. LearnedRewrite (Zhou et al., 2021) uses Monte Carlo Tree Search (MCTS) with learned cost models to explore the space of rule applications, but its performance depends on the accuracy of cost model estimation. LLM-R2 (Li et al., 2024) leverages LLMs for rule selection based on similar query examples. However, assembling its curated demonstration pool entails costly execution of rule combinations per query, and disparate data distributions between demonstration and target queries may lead to suboptimal results.
2.2. Query Hints
Query hints are SQL extensions that provide instructions to the database’s query engine to influence execution plan selection (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021). The following example uses the pg_hint_plan extension (NTT OSS Center DBMS Development and Support Team, 2012) to demonstrate how query hints operate:
HashJoin(employees departments)
SeqScan(employees)
SELECT *
FROM employees e
JOIN departments d ON e.dept_id = d.id
WHERE e.salary > 50000;
As shown above, query hints will provide precise control over the optimizer’s choice of physical operators and access methods: /*+ HashJoin(employees departments) */ compels a hash join for the $employees\bowtie departments$ operation, and /*+ SeqScan(employees)*/ enforces a full table scan of the employees relation, thereby overriding the optimizer’s default cost-based plan selection.
Although existing studies on query hints (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023) have shown promising results, they typically focus on boolean hints applied at the query level (e.g., the enable_nestloop configuration parameter in PostgreSQL), which apply globally and can affect subsequent workload queries. In contrast, pg_hint_plan injects fine‐grained hints at the level of individual joins or cardinality estimates, allowing the use of multiple hint types (e.g., cardinality adjustments) at a more fine-grained level (e.g., specifying join algorithms for particular table pairs) while requiring a more precise selection mechanism.
2.3. LLM and LLM Agents
LLM for Database Tasks. LLMs are increasingly applied to database tasks like text-to-SQL (Gao et al., 2023), SQL workload generation (Lao and Trummer, 2025), data processing (Albalak et al., 2024), and database optimization (Lao et al., 2024a; Zhao et al., 2025; Lao et al., 2024b; Tan et al., 2025; Lao et al., 2025), with recent work focusing on query optimization (Li et al., 2024; Sun et al., 2024; Liu and Mozafari, 2024). Although these methods show LLMs can understand and modify SQL, the inherent complexity of query rewrite presents substantial challenges.
LLM Agents. The primary advantage of LLM agents over LLMs is their ability to collect feedback from the execution environment (e.g., database execution metrics) and dadapt their subsequent reasoning and outputs accordingly. (Wu et al., 2025). They have proven particularly effective in complex problem solving—for example, SWE‐agent in software engineering (Yang et al., 2024) and ChemCrow in chemistry (Bran et al., 2023). QUITE applies this LLM agents methodology to query rewrite by gathering DBMS feedback and adjusting each query candidate accordingly. This performs a feedback-aware refinement, enabling reliable and advanced query rewrite that extends beyond predefined rules or the capabilities of standalone LLMs.
3. System Overview
In this section, we present our system’s end-to-end pipeline (Section 3.1) and describe its core components (Section 3.2).
3.1. System Pipeline
The system pipeline is illustrated in Figure 3. ❶ User provides the SQL query to be optimized along with the corresponding database configuration parameters (e.g., DBMS type, host address). ❷ QUITE obtains metadata (e.g., database statistics, query information) to generate a reasoning chain. This chain, which contains rewrite proposals and corresponding rewritten SQL candidates, is then used to generate the rewritten SQL for downstream enhancement and validation. ❸ QUITE consolidates the SQL candidates generated by the reasoning chains, constructs the prompt and selects and then applies global rewritten SQL enhancements. ❹ QUITE validates the syntactical correctness and semantic equivalence of the rewritten SQL. ❺ QUITE reassesses the rationality of the rewritten SQL to decide whether to proceed or to rewrite the SQL again referring to the rewrite history. In Query Hint Recommender, ❻ QUITE first reduces the hint search space by selecting and creating high-quality hints. Then, ❼ QUITE recommends and injects the most suitable query hints for the rewritten SQL. ❽ Finally, QUITE outputs the final rewritten SQL together with optimization proposals.
3.2. System Components
QUITE has three components: Rewrite Middleware, LLM Agent-Based Query Rewriter, and Query Hint Recommender. They work as follows:
LLM Agent-Based Query Rewriter. The rewriter utilizes a training-free and feedback-aware LLM agent-based Finite State Machine to rewrite queries in multiple stages. First, we formalize query rewrite as a Markov Decision Process (MDP), treating each SQL form as a state, rewrite operations as actions, and cost‐reduction estimates as rewards to guide the LLM’s chain‐of‐thought toward lower‐cost rewrites (see Subsection 5.1). Second, we decompose the rewrite workflow into an FSM with three specialized agents, utilizing a multi‐agent collaboration strategy and constructing an FSM to ensure a stable workflow (see Subsection 5.2).
Rewrite Middleware. The toolkit prepares three specialized tools to enhance the rewrite capability of LLM agents. First, we construct a structured knowledge base using a multi-source query rewrite knowledge integration pipeline to guide the LLM agent to generate high-quality rewrite proposals (see Section 5.1). Second, we develop a hybrid SQL corrector incorporating an LLM agent to validate and correct the syntactical correctness and semantic equivalence of the rewritten SQL (see Section 5.2). Third, we design an agent memory buffer that can extract key information from previous agent conversations, thus reducing agent communication overhead and LLM hallucination in the rewrite process (see Section 5.3).
Query Hint Recommender. The recommender explores the query hints in two steps. First, we carefully construct a promising query hint base, thereby reducing the search space (see Section 6.1). Second, we select appropriate hints for the rewritten query to improve its execution plan (see Section 6.2). Finally, the selected hints are injected into the rewritten SQL to produce the final output.
4. LLM Agent-Based Query Rewriter
In this section, we present our LLM agent–based query rewriter and describe how to construct an efficient, robust query‐rewriting workflow using LLM agents.
Figure 4. Reasoning Chain of MDP-based Reasoning Agent
4.1. MDP-based Reasoning Agent
Query rewrite is fundamentally a sequential decision problem (Finance and Gardarin, 1991b; Zhou et al., 2021), each transformation step changes the query structure and limits future optimization options. Reasoning LLM agents (e.g., DeepSeek-R1 (Guo et al., 2025)), capable of chain-of-thought planning and multi-step logical inference, appear well-suited to this task. However, our experiments show they still produce 13.0% nonequivalent outputs and yield minimal execution improvements (Section 7.2). We observe that query rewrite mirrors a Markov Decision Process, where states are query forms, actions are rewrite steps, and rewards are cost reductions (Puterman, 1990; Liu et al., 2023; Wang, 2025). By framing rewrite in this way, we provide the reasoning LLM agent with a clear, stepwise optimization goal that guides each inference step. Therefore, we design an MDP-based reasoning agent to produce high-performing rewrites.
We model the query rewrite process as a Markov Decision Process $(\mathcal{S},\mathcal{A},\mathcal{T},r,\gamma)$. Starting from the original query $Q$, each state $S_{t}\in\mathcal{S}$ represents the current form of the query with any semantical-level rewrite attempts. At each timestep $t$, the agent chooses either a refinement action $R_{t}$ (e.g., join reordering, predicate pushdown) or a terminal action $\mathcal{A}$ that emits the final rewritten query. This yields the deterministic transition:
$S_{t+1}=\begin{cases}\mathcal{T}(S_{t},R_{t})=S_{t}+R_{t},&\text{if a refinement is applied},\\ \mathcal{T}(S_{t},a)=S_{t}+\mathcal{A},&\text{if rewriting terminates}.\end{cases}$
To guide the search toward efficient rewritten queries, we define a domain‐specific reward function once $S_{t}$ transitions to $S_{t+1}$ under $R_{t}$:
$r(S_{t},R_{t})\;=\;\mathrm{Cost}(S_{t})\;-\;\mathrm{Cost}(S_{t+1}),$
where $\mathrm{Cost}(\cdot)$ denotes the DBMS optimizer’s estimated execution cost (via EXPLAIN) and LLM’s evaluation. Intuitively, each positive reward corresponds to a drop in estimated cost. The agent’s objective is captured by the action‐value function $Q^{\pi}$ under policy $\pi_{\theta}(a\mid S_{t})$, where action $a\in R_{t}$ :
$Q^{\pi}(S_{t},a_{t})\;=\;\mathbb{E}_{\pi}\Bigl{[}\sum_{k=0}^{T-t}\gamma^{k}\,r(S_{t+k},a_{t+k})\;\Bigm{|}\;S_{t},a_{t}\Bigr{]},$
which represents the expected total discounted reward starting from state $S_{t}$ after taking action $a_{t}$. The optimal action‐value function $Q^{*}$ then satisfies the Bellman equation:
$Q^{*}(S_{t},a_{t})\;=\;r(S_{t},a_{t})\;+\;\gamma\,\max_{a^{\prime}\in\mathcal{A}}Q^{*}(S_{t+1},a^{\prime}),$
In practice, we realize the policy $\pi_{\theta}(a\mid S_{t})$ as a reasoning‐specialized LLM (e.g., Deepseek-R1) that generates each action conditioned on the current query state.
As illustrated in Figure 4, the agent transitions from the initial state $S_{0}$ through a sequence of refinement actions $\{R_{0},R_{1},\dots\}$, and finally applies the terminal action $A$ to emit the final rewritten query. Concretely, at each timestep $t$, the agent is prompted with the current SQL form $S_{t}$ and analyzes the query structure via its internal chain‐of‐thought, scores candidate refinements according to the expected cost reduction, and samples the highest‐scoring $R_{t}$.
4.2. FSM-based Query Rewrite
Single-agent approaches often produce suboptimal or invalid rewrites because a single agent is responsible for managing the entire rewrite pipeline (e.g., propose rewrites, verify validity, and assess efficiency), which overloads the model’s context capacity and ultimately increases the risk of hallucination. Inspired by multi-agent frameworks that assign specialized agents to distinct tasks (Hong et al., 2023), we break down query rewrite into discrete stages, each handled by a dedicated LLM agent. To ensure the multi-agent system’s stability and correctness (Han et al., 2024), we implement a Finite State Machine (FSM) to orchestrate stable stage transitions.
4.2.1. Query Rewrite Process Decomposition
Given an environment $\mathcal{E}$ representing a database system, an LLM Agent Group for Query Rewrite is a tuple $\mathcal{A}=(\mathcal{G},\mathcal{M},\mathcal{F},\mathcal{R})$ where:
  • $\mathcal{G}=\{A_{1},...,A_{n}\}$ is a set of cooperative LLM-based agents, each with specialized capabilities $C_{i}$ for tasks.
  • $\mathcal{M}:\mathcal{Q}\times\mathcal{E}\rightarrow\mathcal{S}$ is a transmission that breaks down the input query $Q\in\mathcal{Q}$ into intermediate states $s\in\mathcal{S}$.
  • $\mathcal{F}:\mathcal{S}\times\mathcal{H}\rightarrow\mathcal{Q}^{\prime}$ is a transformation function that produces rewritten queries $Q^{\prime}\in\mathcal{Q}^{\prime}$ based on the current state and historical feedback $h\in\mathcal{H}$ from $\mathcal{E}$.
  • $\mathcal{R}:\mathcal{Q}^{\prime}\times\mathcal{E}\rightarrow\mathbb{R}^{+}$ is a reward function that evaluates rewrite quality through execution metrics and guides iterative refinement.
Figure 5. The Workflow of FSM-based Query Rewrite
Table 1. Specialized Agents in the Finite State Machine
State Agent Agent Role
$S_{1}$ MDP-based Reasoning Agent Generate Reasoning Chain
$S_{1}$ Rewrite Agent Extract & Refine SQL candidates
$S_{2}$ Assistant Agent Validate Syntax & Equivalence
$S_{3}$ Decision Agent Score & Control Termination
Specifically, the Agent takes action $O_{i}$ in state $s_{i}$, the environment provides feedback $O_{i}^{\prime}$, and the Agent updates its state and decides the next action based on the feedback as:
$\mathcal{A}_{s_{i}}\xrightarrow{O_{i}}\mathcal{E}\xrightarrow{O_{i}^{\prime}}\mathcal{A}_{s_{i+1}}$
where $\mathcal{A}_{s_{i}}$ denotes the Agent in state $s_{i}$, and $\mathcal{A}_{s_{i+1}}$ denotes the Agent in state $s_{i+1}$.
In practice, we decompose the rewrite process into four stages: Reasoning, Verification, Decision and Termination, as illustrated in Figure 7. The Reasoning stage generates candidate SQL rewrites. The Verification stage validates syntax and semantic equivalence, correcting any errors identified. The Decision stage determines whether to continue or terminate the rewrite process.
4.2.2. Specialized LLM Agents
As shown in Table 1, we deploy three specialized LLM agents alongside the core MDP-based Reasoning Agent in Section 4.1 to drive the rewrite pipeline for defined FSM stages. Then we propose an LLM agent-based query rewrite methodology detailed in Algorithm 1:
  • Reasoning stage ($S_{1}$): In this stage, the MDP-based Reasoning Agent generates a reasoning chain, containing detailed rewrite proposals and SQL candidates. The Rewrite Agent then serves as a reward model, extracting and refining SQL candidates from the reasoning chain (Lines 3-7).
  • Verification stage ($S_{2}$): In this stage, the Assistant Agent validates syntactical correctness and semantic equivalence by using the hybrid SQL corrector (Lines 8-12).
  • Decision stage ($S_{3}$): In this stage, the Decision Agent reassesses the rewritten SQL’s efficiency with comprehensive report, and determines whether the rewrite process should continue or terminate (Lines 13-19).
1
Input: Original SQL Query $Q_{0}$; LLM Agent $\mathcal{M}$; Tool Repository $T$; Knowledge Base $\mathcal{K}$
Output: Optimized SQL Query $Q^{*}$ with reduced cost
2
3Initialize State $S\leftarrow\textsc{Reasoning}$; Iteration $t\leftarrow 0$; Advanced Knowledge $\mathcal{A}\leftarrow\emptyset$;
4
5while $S\neq\textsc{Termination}\land t\leq T_{\max}$ do
6       if $S=\textsc{Reasoning}$ then
7             $\mathcal{C}\leftarrow\text{Generate-Reasoning-Chain}(\mathcal{M},Q_{0},T)$;
8             $\langle Q^{P},\mathcal{A}\rangle\leftarrow\text{Extract-Rewrite-Proposal}(\mathcal{C})$;
9             $\langle Q^{E},\text{EarlyStop}\rangle\leftarrow\text{Enhance-Query}(\mathcal{M},Q^{P},\mathcal{A})$;
10             $S\leftarrow\text{EarlyStop}\vspace{0.25pt}?\vspace{0.25pt}\textsc{Termination}:\textsc{Verification}$;
11
12       else if $S=\textsc{Verification}$ then
13             while $Q^{E}$ not pass verification do
14                   Correct syntax and equivalence of $Q^{E}$ via $T$;
15
16             end while
17            $S\leftarrow\textsc{Decision}$;
18
19       else if $S=\textsc{Decision}$ then
20             $\mathcal{R}\leftarrow\text{Generate-Report}(Q^{E},Q_{0},\mathcal{M},T)$;
[…]
22                   $Q^{*}\leftarrow Q^{E}$; $S\leftarrow\textsc{Termination}$;
23
24             else
25                   $\mathcal{A}\leftarrow\mathcal{A}\cup\text{Retrieve-Knowledge}(\mathcal{K},Q^{E})$;
26                   $S\leftarrow\textsc{Reasoning}$;
27
28
29       $t\leftarrow t+1$;
30
31 end while
32return $\langle Q^{*},\text{cost}(Q^{*}),\mathcal{A}\rangle$;
Algorithm 1 LLM-Agent-based Query Rewrite Algorithm
Rewrite Agent. After the Reasoning agent generates a chain, the Rewrite agent extracts proposals from each node, groups them into four categories matching our structured knowledge base, evaluates each proposal’s expected reward, and selects the top SQL candidate. Since this candidate may introduce new structural complexities, it cannot be used directly for verification. Therefore, the Rewrite agent performs an SQL Enhancement Process—using the prompt template below—to identify further rewrite opportunities.
[…]
1. Push down predicates or use a CTE structure
2. Simplify calculations and remove redundant predicates
3. Apply other efficiency-preserving transformations
If you think the SQL candidate is well-optimized, just return the original version as the rewritten SQL.
Assistant Agent. The Assistant agent uses the hybrid SQL corrector (Section 5.2) to ensure the syntactically correctness and semantically equivalence of queries. This is performed in two steps:
  • Syntax correction. If syntax errors occur, it attempts up to $K_{\max}$ repairs. If the query remains uncorrected after these attempts, the current FSM iteration is aborted, and control returns to the reasoning stage.
  • Equivalence correction. Once syntax‐correct, the agent submits the query pair $\langle Q_{\text{orig}},Q_{\text{new}}\rangle$ to verify the equivalence and correct nonequivalent queries.
This two-phase process ensures that only valid, semantically equivalent rewrites advance to the Decision stage. Although equivalence verification may return the original format, it avoids the far greater cost of returning nonequivalent SQL while preventing performance degradation from excessive loop iterations.
Decision Agent. The Decision agent implements a cost-aware decision mechanism to determine whether the rewrite process should terminate within budget constraints. It contains two steps:
(i) Generate Comprehensive Report. The Decision agent analyzes the rewritten query across four dimensions, including cost changes, plan characteristics, resource utilization, and other improvements. This multi-dimensional approach overcomes the limitations of relying solely on EXPLAIN cost estimates, preventing the system from accepting misleading or suboptimal rewrites.
(2) Evaluation and Feedback Loop. Based on the generated report, the Decision agent gives a binary judge to decide whether the rewrite process should terminate or not. If True, the Decision agent terminates the FSM process; If False, it retrieves higher-confidence rewrite proposals from the knowledge base, then stores the current query, report, and retrieved knowledge in an agent shared buffer for the next FSM iteration. We deliberately avoid introducing knowledge base retrieval in the initial state, because reasoning models have input content limitations—excessive guidance information can reduce the generation freedom, potentially compromising the quality of reasoning proposal chains (Guo et al., 2025). Finally, the FSM enters the Termination stage and outputs the final rewritten query along with its comprehensive optimization report.
5. Rewrite Middleware
In this section, we introduce the Rewrite Middleware. It comprises three key components: a structured knowledge base of query rewrite strategies (Section 5.1), a hybrid SQL corrector ensuring syntactic validity and semantic equivalence (Section 5.2) and an agent memory buffer for agent context management (Section 5.3).
5.1. Structured Knowledge Base
Preparing the knowledge base involves the following steps: (i) collecting SQL rewrite knowledge from various resources, (ii) filtering untrustworthy content, (iii) enhancing the knowledge, and (iv) retrieving relevant knowledge for specific tasks. The Knowledge Base $\mathcal{K}$ is defined as follows:
Definition 1 (Knowledge Base). A Knowledge Base $\mathcal{K}$ is a set of question and answer (Q&A) pairs, denoted as $\mathcal{K}=\{\langle q_{1},a_{1}\rangle,\langle q_{2},a_{2}\rangle,\allowbreak...,\langle q_{n},a_{n}\rangle\}$. Each Q&A pair $\langle q_{i},a_{i}\rangle$ consists of:
  • A question $q_{i}=\langle Text_{que_{i}},SQL_{que_{i}}\rangle$, where $Text_{que_{i}}$ describes a specific SQL rewrite issue, and $SQL_{que_{i}}$ is the original SQL.
  • An answer $a_{i}=\langle Text_{ans_{i}},SQL_{ans_{i}}\rangle$, where $Text_{ans_{i}}$ explains the strategies used to rewrite $SQL_{que_{i}}$, and $SQL_{ans_{i}}$ is the rewritten SQL query.
Step 1: Collecting rewrite knowledge. Our knowledge is primarily sourced from official database documentation and the database community. Since rewrite strategies are often indirectly presented and scattered throughout official documents (pos, [n.d.]b; mys, [n.d.]; sql, [n.d.]; ora, [n.d.]), we employ LLMs to identify and summarize them into key points. We also collect knowledge from StackOverflow (sta, 2025), as it has abundant real-world query rewrite examples and insights accumulated over decades. We retrieve all Q&A units tagged with “query rewrite” or “query optimization” that include at least one answer. In total, we collected 3,432 valid Q&A units.
Step 2: Filtering untrustworthy content. We begin by filtering out answers that receive more dislikes than likes. Next, we prioritize solutions achieving consensus within the discussion, using an LLM to determine if consensus has been reached. Following this, we employ a majority vote by LLMs to identify the most effective solution from these consensus-based responses. This process yields a total of 241 high-quality Q&A units.
Step 3: Enhancing the knowledge. An LLM first summarizes questions and user’s responses from website Q&A units. These summaries are then augmented with corroborating information from official documents. To do this, we embed user response summaries using Sentence Transformers (Reimers and Gurevych, 2019). These embeddings are then used to query a pre-indexed knowledge base of extracted points, retrieving the top three matches via cosine similarity. Finally, an LLM evaluates these matches to confirm they effectively explain the rewrite strategy underlying the Q&A unit. Suitable matches are incorporated, enriching Q&A units with concise explanations that aid future retrieval.
Step 4: Retrieving relevant knowledge. We classify these units into four categories: join optimization, constant folding, predicate simplification and other, which capture the essential aspects of query rewriting and provides a foundational structure for enabling precise retrieval of information. Since knowledge bases are often context-specific, we use the BM25 algorithm (Robertson et al., 2009) to retrieve the most relevant documents by computing the document relevance.
5.2. Hybrid SQL Corrector
Hallucinations in LLMs can cause them to generate faulty queries, with two primary types of errors: (1) Syntax Error, where a generated SQL violates language grammar and cannot execute; (2) Equivalence Error, where a generated SQL executes but yields results or semantics different from the original SQL. Our proposed hybrid correction method targets both issues.
Syntax Error Correction. Syntax errors, typically minor issues like misspelled keywords or incorrect column/table names, prevent the database from parsing the SQL query. An LLM, leveraging its understanding of SQL syntax and guided by error messages from the DBMS, can efficiently identify and fix these structural flaws. Because such corrections usually involve minimal changes targeting obvious mistakes, the query’s original intended semantics are generally preserved with low overhead during this process.
Equivalence Error Correction. Verifying the equivalence between any pair of SQL queries is known to be an NP-hard challenge (Aho et al., 1979), and whether this problem admits a polynomial-time solution remains an open question. While numerous verification methods (Chu et al., 2017a; Wang et al., 2022, 2024; Ding et al., 2023) rooted in algebraic or symbolic reasoning have been proposed to prove equivalence, they typically operate on constrained query sets (e.g., specific operators, advanced SQL constructs). Recent studies have highlighted the potential of LLMs in deducing query equivalence with high confidence (Singh and Bedathur, 2024; Zhao et al., 2023). However, relying solely on LLMs carries inherent risks. Therefore, we propose a hybrid approach that integrates traditional verification tools with LLM capabilities to achieve broader and more reliable equivalence judgments in two stages.
  • Tool-based Verification. We first use SQLSolver (Ding et al., 2023), an advanced verification tool, for an initial check. Given a pair of queries, $\langle Q_{orig},\allowbreak Q_{rewr}\rangle$, SQLSolver determines their equivalence as either equivalent, nonequivalent, or unknown (due to timeout or other limitations). For query pairs where SQLSolver’s outcome is unknown, verification proceeds to a subsequent stage, utilizing the proposed LLM-based approach.
  • LLM-based Verification and Correction. Leveraging their ability to understand semantic relationships, LLMs can effectively check SQL query equivalence and correct nonequivalent rewrites. However, directly applying LLMs faces challenges, as they may be unsure for some complex queries or incorrectly identify a pair of nonequivalent queries as equivalent. To mitigate these issues, we employ an iterative verification and correction process. In this process, the LLM repeatedly compares the original query $Q_{orig}$ with the rewrite candidate, systematically identifying discrepancies and applying refinements. This iterative loop continues until the LLM confidently establishes query equivalence or a predefined time budget is exhausted. If the time budget is reached without the LLM confidently establishing equivalence, the system defaults to returning the original query ($Q_{rewri}=Q_{orig}$), ensuring that system performance is not compromised by unresolved complex rewrites. Our experiments on diverse workloads show that our method achieves higher rewrite equivalence rates than baselines (Figure 6).
5.3. Agent Memory Buffer
Effective context sharing is crucial in LLM-agent based query rewrite workflows. Simply logging the entire prompt history is problematic: (1) Hallucinations due to redundant context. Lengthy SQL statements and redundant reasoning tokens can distract LLM, causing it to generate incorrect content (Shi et al., 2023). (2) Performance degradation under long-context overhead. Unbounded history growth incurs quadratic attention costs and increases inference latency (Jiang et al., 2023; Kuratov et al., 2024).
To address these, we propose an agent memory buffer $\mathcal{B}$, containing memory slices =$\mathcal{B}=\{m_{1},m_{2},...,m_{n}\}$, where each slice $m_{i}$ is dedicated to a specific category of critical information (e.g., query information, rewrite proposals, retrieval knowledge). By extracting only essential messages across the rewrite process, our buffer curbs hallucinations and bounds context growth, thereby improving both correctness and efficiency in single-agent SQL rewrite workflows.
Table 2. Constructed Hint Base
Query Hint Grammar Description
No Hash Join /*+ NoHashJoin(table table[ table...]) */ Prevents the use of hash joins for specified tables.
No Nested Loop Join /*+ NoNestLoop(table table[ table...]) */ Prevents the use of nested loop joins for specified tables.
No Merge Join /*+ NoMergeJoin(table table[ table...]) */ Prevents the use of merge joins for specified tables.
Row Correction /*+ Rows(a b #10) */ Sets the number of rows for the join result.
NO_MATERIALIZE /*+ NO_MATERIALIZE(table) */ Inlines the CTE to avoid materialization overhead.
6. Query Hint Recommender
Motivation. Query rewrite applies high-level transformations to give the DBMS optimizer a better starting point for plan generation. While this improved starting point is beneficial, the optimizer, which dictates the finer-grained execution operations, can still generate suboptimal plans. This sub-optimality arises because the optimizer’s decisions rely on cost models that are often inaccurate due to outdated statistics and simplifying assumptions (e.g., statistical independence of columns) (Stillger et al., 2001; Marcus et al., 2019; Chen et al., 2023). To exert more fine-grained control over plan selection, we propose embedding query hints (Marcus et al., 2021) within rewritten queries. In this work, we introduce a curated hint base and use LLMs to adaptively select and incorporate effective hints into rewritten queries.
6.1. Hint Base Construction
Most modern database systems (e.g., PostgreSQL, MySQL, Oracle) support query hints, either inherently or by extensions (NTT OSS Center DBMS Development and Support Team, 2012; Oracle Corporation, 2024, 2019). In this work, we use PostgreSQL’s pg_hint_plan extension (NTT OSS Center DBMS Development and Support Team, 2012) to implement our query hint system. Our system supports a broader range of hints (e.g., disabling materialization and setting specific cardinality values), whereas existing works (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021) are limited to Boolean hints.
Selecting Existing Hints. First, we exclude index-related and hardware acceleration hints as they are orthogonal to query rewrite optimization and can introduce hardware-specific dependencies. From the remaining hints, we employ an LLM with the prompt $p_{sel}$ = ”Given a query hint set, you should select the most effective and practical hints for query rewrite and give the reasons.” The LLM’s selections are then further refined based on expert review to ensure an optimal and practical set of hints.
Introducing a Novel CTE Hint. To provide more precise control over Common Table Expression (CTE) execution and address associated performance issues, we introduce a new NO_MATERIALIZE hint. By default, PostgreSQL materializes CTEs (pos, [n.d.]a). This approach can lead to significant performance degradation, particularly when CTEs operate on small datasets or are executed infrequently (El-Helw et al., 2015; Zhou et al., 2007), due to the overhead of writing and reading intermediate results. Furthermore, the decision to materialize a CTE, as opposed to inlining it, directly impacts the optimizer’s ability to perform effective predicate pushdown into subqueries, especially after query rewrite. Our NO_MATERIALIZE hint directs the optimizer to inline the targeted CTE, expanding its definition into the main query as if it were a FROM-clause subquery. This hint enhances query performance by eliminating materialization overhead and enabling more effective plan optimizations, such as improved predicate pushdown. Ultimately, these selected and newly introduced hints constitute our query hint base $\mathcal{H}$, detailed in Table 2.
6.2. Hint Injection
This section details our approach to using LLMs for identifying operations that could benefit from hints and then recommending appropriate ones. Unlike existing methods (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021) that recommend hints at the query level (based on global parameters like GUCs), where a single hint applies to all operations within that query, our strategy targets a more granular level. We found that query-level hints can be suboptimal, as a hint beneficial for one operation might negatively impact others. Therefore, we apply hints at a finer-grained level, such as to individual tables or specific operations within a query, to maximize positive impacts.
First, given the physical plan of $Q^{\prime}$, database semantics and data distribution statistics, we identify potential improvements that can be achieved through hints: (1) For every cardinality estimation in the plan, an LLM judges its reasonableness. If an estimation is deemed unreasonable, the LLM provides a recommended range and justification. (2) For all join operations, the LLM assesses whether the selected join operator (such as hash, merge, or nested-loop join) is suitable given the available statistics. Second, an LLM is then invoked to construct the most effective hints based on the obtained analysis and the query plan. It recommends join methods for specific tables and rectifies cardinality estimations. Regarding NO_MATERIALIZE, we apply this hint only if a CTE appears exactly once or processes a small, simple dataset. Finally, the proposed hints are then injected into $Q^{\prime}$, producing the final optimized query $Q^{*}$.
Our LLM-guided hint injection, unlike general existing methods, pinpoints costly operations and generates finer-grained hint recommendations. This enhances the explainability of the recommendations and leads to further improvement. Moreover, in contrast to ML-based hint recommendation approaches (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021) that rely on pre-trained models and multi-plan enumeration, our method requires zero training yet surpasses their performance.
7. Experiment
7.1. Experimental Set up
Testbed. Our experiments were conducted on a cloud server featuring a 32-core Intel Xeon Platinum 8352V CPU, 251 GB of RAM, a 942 GB SSD, and PostgreSQL v14.13. We utilize a NVIDIA A100-SXM4 GPU with 40 GB of memory to run the rewrite process of LLM-R2.
Workloads. We employ three widely recognized datasets for our evaluations. (1) TPC-H (TPC, nd) is a well-known OLAP benchmark containing 62 columns and 22 query templates. We generate 63 queries from these templates, excluding Q15 due to its use of CREATE VIEW, a format unsupported by the rewrite engines of two baseline methods (Zhou et al., 2021; Li et al., 2024). We used a scale factor (SF) of 10. (2) DSB (Ding et al., 2021) is adapted from the TPC-DS benchmark (Nambiar and Poess, 2006) and is notable for its complex data distribution and challenging long-context query templates, which consists of 52 query templates. We generate 156 queries based on these templates with $SF=10$. (3) Calcite (apa, 2021) is a real-world workload used to evaluate rewrite rules. We randomly select 58 queries and generate 10G of data with uniform distributions following previous work (Dong et al., 2023).
Baselines. We compare our system against the following state-of-the-art query rewrite baselines:
  • LearnedRewrite (LR) (Zhou et al., 2021) uses a Monte Carlo Tree Search (MCTS) algorithm with a learned cost estimation model to explore a SQL policy tree of rewrite rule orders. We adopt the original implementation and use the cost estimation model provided in their official code repository.
  • LLM-R2 (Li et al., 2024) is a rule-based query rewrite system built on Apache Calcite (Begoli et al., 2018). It uses LLM’s In-Context Learning (ICL) ability to select query rewrite rules according to high quality demonstration queries.
  • LLM Agent rewrites queries using an individual LLM agent. To ensure a fair comparison, this LLM agent can access target database statistics and our SQL corrector for self-reflection.
  • R-bot (Sun et al., 2024) is an LLM-based system that leverages an embedded retrieval-augmented generation (RAG) knowledge base and reflective reasoning to select promising rewrite rules.
Table 3. Different Methods on Query Latency (s)
Methods TPC-H (SF=10) DSB (SF=10) Calcite (SF=10)
Mean Median 75th 95th Mean Median 75th 95th Mean Median 75th 95th
Original 69.84 9.64 32.75 300.00 32.62 4.85 10.29 300.00 23.88 2.58 8.64 122.36
LearnedRewrite 37.57 9.97 30.55 202.31 31.93 5.14 15.18 251.42 22.88 2.32 8.09 122.58
LLM-R2(GPT-4o) 57.09 9.30 30.50 300.00 10.53 4.11 8.36 32.88 16.98 2.11 7.75 50.97
LLM-R2(Claude-3.7) 57.68 9.89 30.07 300.00 9.11 3.82 8.24 27.48 16.96 2.55 8.19 50.07
LLM-R2(DS-R1) 56.73 9.25 29.33 300.00 9.16 3.84 8.21 24.21 16.90 2.05 7.71 49.91
LLM Agent (GPT-4o) 69.03 10.00 27.02 300.00 33.44 4.39 9.65 300.00 15.47 2.03 7.32 35.34
LLM Agent (DS-R1) 42.79 9.94 27.95 300.00 13.53 3.87 7.93 30.18 15.18 1.65 7.55 36.70
LLM Agent (DS-V3) 60.12 9.22 29.92 300.00 32.34 4.19 9.99 300.00 16.49 2.53 7.72 45.15
LLM Agent (Claude-3.7) 57.46 11.04 36.99 300.00 25.32 4.38 8.84 300.00 15.67 1.59 7.68 44.54
R-Bot (GPT-4o) 32.64 9.48 26.25 172.20 21.75 4.26 8.80 119.19 22.21 2.34 8.35 122.14
R-Bot (Claude-3.7) 33.89 9.71 29.32 195.51 19.89 4.27 8.99 112.24 23.38 2.60 8.64 122.95
R-Bot (DS-R1) 33.70 9.88 28.38 176.90 21.21 4.44 9.11 113.91 22.71 2.52 8.43 122.89
[…]
(a) TPC-H Benchmark
(b) DSB Benchmark
(c) Calcite Benchmark
Figure 6. Rewrite Equivalence and Improvement Numbers on Different Benchmarks
QUITE Settings. We use DeepSeek-R1 as the reasoning agent due to its strong analysis capability, and Claude-3.7-Sonnet as the rewrite, decision and assistant agents as default. A temperature of 0 is applied to all LLMs to ensure output stability  (DeepSeek-AI, 2025; Anthropic, 2025). The number of iterations is limited to 2 to prevent excessive FSM rewrite cycles. All LLM-involved experiments employ the identical temperature configurations used in our approach.
Evaluation Metrics. (1) Query Execution Latency. The duration required to complete a query, reported using the average, median, and 95th percentile latency values. (2) Rewritten Equivalence Rate. The fraction of rewritten queries that produce results matching those of the original queries. In our work, equivalence is determined by comparing the execution outputs of the original and rewritten queries. (3) Rewritten Improvement Rate. The percentage of rewritten SQL queries that show a performance improvement. A performance gain is considered significant if it results in at least a 10% reduction in execution time.
Query Evaluation Approach. Before each query, we restart the evaluation database to clear all caches. The query execution process involves an initial warm-up run (to mitigate cold start impacts), followed by three measured executions. We report the average execution time of these three runs. Queries that run for more than 300 seconds are terminated and their time is recorded as 300 seconds (Li et al., 2024), to reflect impractical execution durations. If a rewritten query produces different output than the original or contains syntax errors, its equivalence metric is marked as False. To simulate a realistic query optimization scenario when dealing with such nonequivalent queries, we assign the rewritten query the execution time of its original counterpart.
7.2. Performance Comparison
Query Execution Latency. As presented in Table 3, queries optimized by QUITE demonstrate significantly reduced execution times, consistently outperforming all baseline methods. Across the TPC-H, DSB, and Calcite benchmarks, QUITE achieves substantial average execution time reductions. Specifically, QUITE reduces execution time by 31.9%, 81.7%, and 56.4% compared to LR; 54.9%, 35.8%, and 41.0% over LLM-R2; and 40.2%, 56.8%, and 34.3% against LLM Agent. When evaluated against R-Bot, QUITE achieves reductions of 21.6%, 70.6%, and 55.1%. QUITE’s advantages are further illustrated by the rewritten improvement rates shown in Figure 6, where QUITE achieves the highest ratios of 31.7%, 37.8%, and 34.5% on TPC-H, DSB, and Calcite, respectively, indicating strong generalization capabilities across different workloads.
The optimization benefits of QUITE over baselines are more pronounced on complex datasets. Notably, the improvement margin of QUITE over LR is substantially larger for the DSB dataset compared to the TPC-H and Calcite datasets. This disparity arises primarily because: (1) The LR we use is trained on the TPC-H, LR can proficiently identify effective rewrite rules and navigate their constrained search spaces. This existing proficiency limits the remaining optimization potential, thereby reducing the margin for QUITE to demonstrate substantial further improvements. (2) LR employs a greedy search within its MCTS algorithm, which proves less effective in the expansive search spaces of complex datasets like DSB, unlike its performance on simpler datasets (e.g., Calcite) with smaller search spaces and inherently less optimization potential.
Figure 7. Ablation Study of Hint Injection on DSB Benchmark
Rewritten Equivalence. Figure 6 details the number of equivalent and improved query rewrites produced by each method across the datasets. Notably, direct rewrites from the LLM Agent exhibit significant semantic equivalence issues. In contrast, QUITE achieves substantially higher equivalence rates–attaining 100%, 96.8%, and 98.3% on TPC-H, DSB, and Calcite, respectively. This superior performance is attributable to the multi-stage correction mechanisms integrated into QUITE’s iterative FSM-based process.
Figure 8. Ablation Study of Hybrid SQL Corrector
7.3. Ablation Study
We design our ablation study from four perspectives by selectively replacing or removing system components.
Rewrite Middleware. To assess the hybrid SQL corrector, we removed it and evaluated its impact on query equivalence and execution efficiency using the DSB dataset. As shown in Figure 8, judge accuracy measures the proportion of correct executions among all decisions made by the tool or LLM while coverage ratio indicates the proportion of queries for which the tool or LLM was able to make a definitive decision.Its removal significantly degrades performance: nonequivalent queries increase from 5 to 11, unchanged queries by 13, and improved queries drop by 17. These results highlight the corrector’s essential role in ensuring query quality and efficiency.
We also evaluate the impact of the structured knowledge component. As shown in Table 4, QUITE consistently outperforms its variant without this module across all metrics. The improvement stems from high-quality Q&A pairs and domain knowledge embedded in the knowledge base, which effectively guide more optimized rewrites.
Table 4. Ablation Study of Structured Knowledge Base
Methods DSB (SF=10)
Mean Median 75th 95th
Without Knowledge 9.96 3.94 8.49 24.81
With Knowledge 6.08 3.78 7.98 21.43
Table 5. Ablation Study of the Reasoning Agent and FSM
Methods DSB (SF=10)
MDP-based Reasoning Agent FSM Mean Median 75th 95th
13.48 4.84 10.22 33.64
12.72 4.80 9.93 29.91
6.08 3.78 7.98 21.43
LLM-Agent-based Query Rewrite FSM To evaluate the impact of the FSM framework within our system, we performed two modifications: (1) the loop responsible for equivalence checking was removed, and (2) the decision stage was eliminated. These changes transformed the process into a linear, end-to-end pipeline while retaining all components and agent tools involved in query rewriting. At the same time, to further assess the impact of the MDP-based decision-making on the Reasoning agent’s capabilities, we conducted an ablation study by removing the MDP component. As shown in Table 5, QUITE outperforms its variant across all evaluation metrics. Notably, the average execution time with the MDP is reduced by 52.2% compared to the version without it, demonstrating that the MDP significantly enhances the effectiveness of the Reasoning agent.
Table 6. Ablation Study of QUITE using Different LLMs
Methods DSB (SF=10)
Mean Median 75th 95th
Deepseek-R1 + GPT-4o 31.38 4.64 9.56 300.00
Deepseek-R1 + Deepseek-V3 28.07 4.43 9.07 300.00
Deepseek-R1 + Claude-3.7 6.08 3.78 7.98 21.43
Table 7. Ablation Study of Hint Construction Methods
Methods DSB (SF=10)
Mean Median 75th 95th
QUITE (without hints) 6.08 3.78 7.98 21.43
QUITE (with Bao’s Hint Base) 6.93 4.65 8.46 20.41
QUITE (with Our Hint Base) 5.85 3.35 7.63 19.97
Table 8. Robustness Study of Different Data Sizes
Methods TPC-H (SF=1) TPC-H (SF=10) TPC-H (SF=30)
Mean Median 75th 95th Mean Median 75th 95th Mean Median 75th 95th
Original 29.48 0.69 1.20 300.00 69.31 9.42 30.08 300.00 64.76 20.61 46.69 300.00
LearnedRewrite 1.46 0.52 1.10 3.91 37.57 9.97 30.55 202.31 52.59 18.90 43.50 300.00
LLM-R2(Deepseek-R1) 15.38 0.69 1.22 7.69 56.73 9.25 29.33 300.00 54.67 18.30 58.11 300.00
R-Bot (GPT-4o) 1.49 0.66 1.03 6.38 32.64 9.48 26.25 172.20 54.54 19.85 61.12 300.00
QUITE 1.02 0.64 1.01 3.79 26.06 9.28 24.19 75.64 40.65 16.89 37.96 131.97
Table 9. Data-Aware Analysis on TPC-H Benchmark
Methods TPC-H (SF=1) TPC-H (SF=10) TPC-H (SF=30)
Mean Median 75th 95th Mean Median 75th 95th Mean Median 75th 95th
Original 29.48 0.69 1.20 300.00 69.31 9.42 30.08 300.00 64.76 20.61 46.69 300.00
QUITE (Rewrite on SF=1) 1.02 0.64 1.01 3.79 36.87 13.53 29.02 279.51 54.36 25.02 43.03 289.48
QUITE (Rewrite on SF=10) 1.08 0.68 1.22 3.79 26.06 9.28 24.19 75.64 45.90 17.44 57.26 155.72
QUITE (Rewrite on SF=30) 5.72 0.50 1.17 5.17 30.50 9.96 24.29 96.68 40.65 16.89 37.96 131.97
Query Hint Injection We evaluate the impact of hint injection on queries rewritten by QUITE. As shown in Table 3, the insertion of hints leads to a further decrease in average execution times by 1.8%, 3.8%, and 0.3% on TPC-H, DSB, and Calcite, respectively. These performance gains demonstrate that appropriate hints enable additional fine-grained optimization, even when queries have already been extensively optimized through rewriting. To highlight the advantage of our LLM-based hint injection approach, we compare its performance on rewritten queries against using LLMs to recommend Bao’s candidate hint sets. As shown in Table 7 and Figure 7, within the hint space defined at the GUC level provided by Bao, the overall trend indicates degraded performance.
Furthermore, Figure 7 presents a comparison of execution times for 24 positively optimized queries on DSB, both before and after hint injection, as well as against the original queries. Notably, some queries that exhibited degraded performance after rewriting were able to exceed their original execution efficiency through proper hint injection. This demonstrates that hint injection can effectively recover and enhance performance in cases where the initial rewrites were suboptimal.
Base Model Selection.To evaluate the role of the base models within our framework, we conducted experiments using different model combinations. DeepSeek-R1 was retained as the Reasoning agent due to its low cost and strong chain-of-thought reasoning capabilities, making it indispensable for this role. For the Decision agent, we tested three alternatives: Claude-3.7-Sonnet, DeepSeek-V3, and GPT-4o. As shown in Table 6, rewrites using Claude-3.7-Sonnet as the Decision agent significantly outperform those from the other two models. This is due to its strengths in summarization and decision-making, which require long-context understanding, precise reasoning, and accuracy—areas where Claude-3.7-Sonnet excels. In contrast, DeepSeek-V3 shows higher hallucination rates, and GPT-4o offers lower accuracy and weaker long-context handling, making both less suitable for our framework.
Table 10. Cost Analysis on TPC-H Benchmark
Methods Average Time (s) Cost (\$)
LLM-R2 (Claude-3.7) 399.30 0.93
R-Bot (Claude-3.7) 256.96 20.94
QUITE (DS-R1 + Claude-3.7) 291.59 6.76
7.4. Robustness Study
We assess the robustness of QUITE by testing stability across data scales. We scale the TPC-H dataset from SF = 10 to SF = 1 and SF = 30, then execute rewritten queries on all three versions. Table 8 shows that QUITE delivers significant optimizations at each scale and outperforms all baselines. In contrast, LearnedRewrite and LLM-R2 often produce rewrites that worsen performance for complex queries.
We evaluate the robustness of the QUITE framework by analyzing its stability. The stability assessment demonstrates that QUITE generalizes effectively across diverse data scales. To evaluate the stability of QUITE across varying database sizes, we extended the primary TPC-H $SF=10$ dataset by generating proportionally scaled versions at scale factors of $SF=1$ and $SF=30$. For each evaluated method, queries rewritten based on the $SF=10$ dataset were executed and timed on the $SF=1$, $SF=10$, and $SF=30$ datasets. As shown in Table 8, QUITE consistently achieved substantial optimization improvements across all scales, outperforming all baseline methods.
7.5. Further Analysis
Data-Aware Analysis. We evaluated the sensitivity of QUITE using a cross-testing strategy. Specifically, QUITE generated three sets of rewritten queries, each optimized for the data statistics of a specific TPC-H scale: $SF$=1, $SF$=10, and $SF$=30. Each set was subsequently executed across all three datasets. As shown in Table 9, the best performance consistently occurred when queries were executed on the dataset for which they were originally optimized. This result confirms that QUITE effectively distinguishes between different data distributions and produces rewrites that are well-adapted to the target datasets.
Cost Analysis. As shown in Table 10, LLM-R2 incurs high time expenditure due to demonstration construction, while R-Bot consumes substantial tokens during stepwise rewriting. In contrast, QUITE offers a balanced trade-off between time expenditure and cost. By generating much of its output with DeepSeek-R1, it greatly reduces monetary cost. Additionally, it avoids pre-training and extensive knowledge base construction. To further control execution time, we cap FSM iterations, ensuring reliable output quality without incurring excessive time expenditure.
8. Conclusion
This paper presents QUITE, a training-free and feedback-aware system that smartly leverages LLMs to rewrite SQL queries into semantically equivalent forms with significantly improved performance. We rethink the query rewrite process and identify unique opportunities to support a wider range of query patterns and rewrite strategies with good scalability. The implemented QUITE  system significantly outperforms state-of-the-art methods in both query performance and query coverage on widely used benchmarks.

### v2

@@
 Abstract.
-Query rewrite transforms SQL queries into semantically equivalent forms that run more efficiently.
-Existing approaches mainly rely on predefined rewrite rules, but they handle a limited subset of queries and can cause performance regressions.
-This limitation stems from three challenges of rule-based query rewrite: (1) it is hard to discover and verify new rules, (2) fixed rewrite rules do not generalize to new query patterns, and (3) some rewrite techniques cannot be expressed as fixed rules.
+Query rewrite transforms a given SQL query into a semantically equivalent query that can be executed more efficiently.
+Existing approaches mainly rely on predefined rewrite rules.
+However, they can only handle a small subset of queries and may lead to performance regressions.
+This limitation arises from the intrinsic challenges of rule-based query rewrite: (1) it is hard to discover and verify new rewrite rules, (2) fixed rewrite rules do not generalize to new query patterns and are insufficient for handling complex queries, and (3) some rewrite techniques cannot be expressed by fixed rules.
 Motivated by the fact that human experts exhibit significantly better rewrite ability but suffer from scalability, and Large Language Models (LLMs) have demonstrated nearly human-level semantic and reasoning abilities, we propose a new approach of using LLMs to rewrite SQL queries beyond rules.
@@
 Secondly, we develop a rewrite middleware to enhance the ability of LLMs to generate optimized query equivalents.
-Finally, we employ a novel hint injection technique to improve execution plans for rewritten queries.
-Extensive experiments show that QUITE reduces query execution time by up to 35.8 % over state-of-the-art approaches and produces 24.1 % more rewrites than prior methods, covering query cases that earlier systems did not handle.
-PVLDB Reference Format:
-PVLDB, 14(1): XXX-XXX, 2020.
-This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
-Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
-For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
-Copyright is held by the owner/author(s).
-Publication rights licensed to the VLDB Endowment.
-Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
-doi:XX.XX/XXX.XX
+It includes a structured knowledge base for domain knowledge, an SQL corrector to ensure equivalence, and an agent memory buffer to manage critical rewrite contexts.
+Finally, we employ a novel hint injection technique to produce better execution plans for rewritten queries.
+Extensive experiments show that our method achieves up to a 35.8% reduction in query execution over state-of-the-art approaches.
+Moreover, it delivers 24.1% additional rewrites compared to prior methods, which cover query cases that earlier approaches could not handle.
 1. Introduction
-Inefficient SQL queries remain a long-standing challenge in many database applications.
-These inefficiencies often arise from poorly written queries, which are typically produced by inexperienced users or automatically generated by systems (Li, 2019; Zhou et al., 2020; Pirahesh et al., 1992).
-Query rewrite is a promising approach that transforms an original SQL query into a semantically equivalent form with significantly better performance (Finance and Gardarin, 1991a; Pirahesh et al., 1992; Wang et al., 2022).
-An effective rewrite can improve the execution time of a poorly written SQL query by several orders of magnitude (Wang et al., 2022; Zhou et al., 2021).
+Inefficient SQL queries remain a long-standing challenge in many database applications, often arising from poorly written queries by inexperienced users or automatically generated systems (Li, 2019; Zhou et al., 2020; Pirahesh et al., 1992).
+Query rewrite is a promising approach that transforms an original SQL into a semantically equivalent form with better performance (Finance and Gardarin, 1991a; Pirahesh et al., 1992; Wang et al., 2022).
+An effective rewritten query can improve the execution time of a poorly written SQL query by several orders of magnitude (Wang et al., 2022; Zhou et al., 2021).
 Existing query rewrite approaches typically rely on predefined rewrite rules (Begoli et al., 2018; Graefe and McKenna, 1993; PostgreSQL Global Development Group, 2024; Zhou et al., 2021, 2023; Bai et al., 2023; Li et al., 2024; Sun et al., 2024).
@@
 Recently, some approaches have explored the use of Large Language Models (LLMs) to select rewrite rules and determine their application order based on domain knowledge (Li et al., 2024; Sun et al., 2024).
-Figure 1.
-An Example of Using Strategies beyond Rewrite Rules and Using Context-Aware Analysis.
-(b) boosts speed 37.5x over (a) by using a common table expression (CTE) to eliminate redundant calculations.
-(c) leverages data characteristics by joining sale (10 M rows) with prod (100 K rows) yields a 100 K-row intermediate (vs.
-1.1 M rows in (b) ), achieving a 2.79× speedup.
-Figure 2. An Example of Using Query Intent
-However, even state-of-the-art query rewrite approaches can handle only a small subset of queries and may lead to performance regressions (Zhou et al., 2021; Li et al., 2024; Sun et al., 2024).
+Figure 1. An Example of Using Strategies Beyond Rewrite Rules (Example 1.1) and Using Context-Aware Analysis (Example 1.2)
+Figure 2. An Example of Using Query Intent (Example 1.3)
+However, the above approaches can handle only a small subset of queries and may lead to performance regressions (Zhou et al., 2021; Li et al., 2024; Sun et al., 2024).
 This limitation stems from the intrinsic challenges of rule-based query rewrite.
@@
 (1) Humans can perform rewrites that follow rules but go beyond the expressiveness of pattern matching (Example 1.1).
+(2) Humans can conduct context-aware rewrites by considering the underlying data distributions and cost estimates provided by the query optimizer (Example 1.2).
 (3) Experts can apply intent-oriented rewrites that capture the underlying goal of the query, which are not easily expressed as explicit rules (Example 1.3).
-(3) Humans can conduct context-aware rewrites by considering the underlying data distributions and cost estimates provided by the query optimizer (Example 1.3).
 Example 0.
@@
 Example 0.
+Context-Aware Query Rewrite.
+In Figure 1 (c), instead of directly joining two large tables (the sale with 10 M rows and the ttime with 1.4 M rows), we first joins the sale table with the prod table (100 K rows).
+This highly selective join would produce an intermediate result of only  100 K rows.
+Next we join this with ttime table and yield a 2.79 $\times$ speedup over (b) in our experimental setup.
+Such rewrites, which reduce intermediate size, are rare in rule-based systems and require analyzing data distribution and selectivity.
+Example 0.
 Query Rewrite based on Query Intent.
 Figure 2 (a) shows a real-world query from Apache Calcite (apa, 2021).
-Rule-based rewrite cannot optimize this query (Figure 2 (b).
+Rule-based rewrite cannot optimize this query (Figure 2 (b)).
 Identifying that the query intent is to sum salaries directly from the EMP table, human can simplify the original query by eliminating unnecessary joins and sub-queries that do not contribute to the intent of the query, achieving a 2.17 $\times$ execution improvement in our experimental setup.
-Example 0.
-Context-Aware Query Rewrite.
-In figure 1 (c), instead of directly joining two large tables (the sale with 10 M rows and the ttime with 1.4 M rows), the query first joins the sale table with the prod table (100 K rows).
-This highly selective join would produce an intermediate result of only  100 K rows.
-Next we join this with ttime table and this yields a 2.79 $\times$ speedup over (b) in our experimental setup.
-Such rewrites, which reduce intermediate result size, are rare in rule-based systems and require analyzing data distribution and selectivity.
 Although human experts can produce significantly better SQL rewrites than rule-based methods, manual rewrite suffers from poor scalability, especially when there are thousands or even more poorly written queries to optimize in cloud environments.
@@
 First, to enable LLMs to capture query data characteristics and leverage database feedback, we design a multi-agent framework controlled by a finite state machine (FSM) to iteratively refine the rewrite process.
-We decompose the complex rewrite process into a set of subtasks, each handled by a specialized LLM agent to reduce hallucinations and improve rewrite ability.
-Second, to enhance the ability of LLMs to generate high-quality query rewrites, we provide a rewrite middleware that includes: (1) a structured knowledge base extracted from official documentation and query rewrite discussions on web forums, (2) a hybrid SQL corrector to ensure query equivalence, and (3) an agent memory buffer that manages important rewrite contexts.
+We decompose the complex rewrite process into a set of subtasks, each handled by a specialized LLM agent to reduce hallucinations and improve rewrite ability (addressing C1, C2).
+Second, to enhance the ability of LLMs to generate high-quality query rewrites, we provide a rewrite middleware that includes: (1) a structured knowledge base extracted from official documentation and query rewrite discussions on web forums, (2) a hybrid SQL corrector to ensure query equivalence, and (3) an agent memory buffer that manages important rewrite contexts (addressing C1, C2).
 Third, we employ a novel hint injection technique to generate better execution plans for rewritten queries.
 This ensures that query execution exactly follows the optimized plan recommended by LLM agents, avoiding unintended modifications by the query optimizer due to inaccurate cost estimations.
-Additionally, this technique enables further improvements by incorporating physical-level optimizations.
+Additionally, this technique enables further improvements by incorporating physical-level optimizations (addressing C2).
 We extensively evaluate QUITE against state-of-the-art query rewrite systems on widely used benchmarks (e.g., TPC-H, DSB, and Calcite) and show that our system delivers up to a 21.9% speedup in query execution performance over the best-performing alternative.
@@
   • We propose QUITE, a training-free and feedback-aware query rewrite system that leverages LLM agents to support a broader range of query patterns and rewrite strategies than existing methods (Section 3).
+  • We design a multi-agent framework controlled by a finite state machine to enable LLMs to use external tools and enhance the rewrite process with database feedback (Section 4).
   • We develop a rewrite middleware including a structured knowledge base, a hybrid SQL corrector, and an agent memory buffer to enhance the rewrite capability of LLM agents (Section 5).
-  • We design a multi-agent framework controlled by a finite state machine to enable LLMs to use external tools and enhance the rewrite process with database feedback (Section 4).
   • We propose a hint injection technique to produce better execution plans for rewritten queries. (Section 6)
@@
 Query rewrite is a process of transforming a query into another query that is semantically equivalent but with improved performance.
-It is a distinct step prior to query optimization, operating at the level of the user-accessible application programming interface (API) (e.g., SQL).
+It is a preliminary step prior to query optimization, operating at the level of the user-accessible application programming interface (API) (e.g., SQL).
 Definition 2.2 (Query Equivalence).
 Two queries are considered equivalent if they produce the same result for any valid instance of the database schema.
-Conventional studies treat query rewriting as a transformation between plain SQL statements, typically using rule-based methods. There are two research directions:
+Conventional studies treat query rewrite as a transformation between plain SQL statements, typically using rule-based methods. There are two lines of work:
 (1) Discovery of new rules.
-Approaches such as WeTune (Wang et al., 2022) automatically generate and verify logical plan transformations, but they are limited to certain operator types or algebraic equivalences.
-Techniques based on equality saturation (Nandi et al., 2021) and user-guided tools like QueryBooster (Bai et al., 2023) lack full autonomy and broad applicability.
-More recently, GENRewrite (Liu and Mozafari, 2024) uses LLMs to suggest rules in natural language.
-Yet GENRewrite must overcome challenges in formally verifying equivalence, preventing incorrect suggestions, and choosing the best candidates.
+Approaches such as WeTune (Wang et al., 2022) automatically generate and verify logical plan transformations, but they are limited to specific operator types and algebraic equivalences.
+Equality saturation (Nandi et al., 2021) and user-guided tools like QueryBooster (Bai et al., 2023) lack full autonomy and broad applicability.
+Recently, GenRewrite (Liu and Mozafari, 2024) uses LLMs to suggest rules in natural language.
+However, it still struggles with optimal rewrite selection because executing all candidates to measure latency incurs prohibitive overhead and its off-the-shelf verifiers permit equivalence checking for only a subset of rewrite pairs (Liu and Mozafari, 2024; Dong et al., 2023).
 (2) Effective use of existing rules.
@@
 LearnedRewrite (Zhou et al., 2021) uses Monte Carlo Tree Search (MCTS) with learned cost models to explore the space of rule applications, but its performance depends on the accuracy of cost model estimation.
-LLM-R2 (Li et al., 2024) leverages LLMs for rule selection based on similar query examples.
-However, assembling its curated demonstration pool entails costly execution of rule combinations per query, and disparate data distributions between demonstration and target queries may lead to suboptimal results.
+LLM-R2 (Li et al., 2024) uses a trained selection model to retrieve the past rewrite query and rule demonstrations similar to the input SQL, and leverages them as an in-context prompt for the LLM to select promising rewrite rules.
+However, it still incurs high costs by executing every query demonstration on each rule set and faces distribution shifts between the demonstration and input SQL, since structurally similar queries may differ in row count, leading to suboptimal rule selection.
 2.2. Query Hints
-Query hints are SQL extensions that provide instructions to the database’s query engine to influence execution plan selection (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021).
-The following example uses the pg_hint_plan extension (NTT OSS Center DBMS Development and Support Team, 2012) to demonstrate how query hints operate:
+Query hints are SQL extensions that provide instructions to the database’s query engine to influence the selection of execution plans (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021).
+These hints allow users to exert precise control over how the database optimizer chooses physical operators and access methods.
+As shown below for pg_hint_plan extension (NTT OSS Center DBMS Development and Support Team, 2012), the hint /*+ HashJoin(employees departments) */ forces the query engine to use a hash join for the $employees\bowtie departments$ operation, while the /*+ SeqScan(employees)*/ enforces a full table scan of the employees relation.
+These hints can override the optimizer’s default cost-based plan selection.
 HashJoin(employees departments)
@@
 WHERE e.salary > 50000;
-As shown above, query hints will provide precise control over the optimizer’s choice of physical operators and access methods: /*+ HashJoin(employees departments) */ compels a hash join for the $employees\bowtie departments$ operation, and /*+ SeqScan(employees)*/ enforces a full table scan of the employees relation, thereby overriding the optimizer’s default cost-based plan selection.
 Although existing studies on query hints (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023) have shown promising results, they typically focus on boolean hints applied at the query level (e.g., the enable_nestloop configuration parameter in PostgreSQL), which apply globally and can affect subsequent workload queries.
@@
 LLMs are increasingly applied to database tasks like text-to-SQL (Gao et al., 2023), SQL workload generation (Lao and Trummer, 2025), data processing (Albalak et al., 2024), and database optimization (Lao et al., 2024a; Zhao et al., 2025; Lao et al., 2024b; Tan et al., 2025; Lao et al., 2025), with recent work focusing on query optimization (Li et al., 2024; Sun et al., 2024; Liu and Mozafari, 2024).
-Although these methods show LLMs can understand and modify SQL, the inherent complexity of query rewrite presents substantial challenges.
+Although these methods show that LLMs can understand and modify SQL, the inherent complexity of query rewrite presents substantial challenges.
 LLM Agents.
-The primary advantage of LLM agents over LLMs is their ability to collect feedback from the execution environment (e.g., database execution metrics) and dadapt their subsequent reasoning and outputs accordingly.
-(Wu et al., 2025).
+The primary advantage of LLM agents over LLMs is their ability to collect feedback from the execution environment (e.g., database execution metrics) and dadapt their subsequent reasoning and outputs accordingly (Wu et al., 2025).
 They have proven particularly effective in complex problem solving—for example, SWE‐agent in software engineering (Yang et al., 2024) and ChemCrow in chemistry (Bran et al., 2023).
 QUITE applies this LLM agents methodology to query rewrite by gathering DBMS feedback and adjusting each query candidate accordingly.
-This performs a feedback-aware refinement, enabling reliable and advanced query rewrite that extends beyond predefined rules or the capabilities of standalone LLMs.
+This performs a feedback-driven refinement, enabling reliable and advanced query rewrite that extends beyond predefined rules or the capabilities of standalone LLMs.
+2.4. Comparison with Other Works
+Instead of focusing on finding optimal rewrite rules with limited coverage (Zhou et al., 2021; Li et al., 2024; Sun et al., 2024), QUITE addresses the core aspects of query rewriting by identifying semantic and physical rewrite bottlenecks through real-time interaction between LLM agents and the database, enhancing the detection of performance issues in the input SQL.
+Furthermore, QUITE generates a reasoning chain with natural language rewrite proposals and rewrites the SQL through multi-agent collaboration within a carefully designed Finite State Machine (FSM) pipeline and agent tools, ensuring equivalence and effectiveness.
+This approach enhances adaptability and efficiency in context-aware rewrites, compared to rule-based systems that miss complex query patterns and context-specific optimizations.
+In contrast to GenRewrite (Liu and Mozafari, 2024), which directly utilizes LLMs to generate rewrite rules for rewriting, QUITE considers key factors such as the SQL’s physical plan, the database distribution, and the domain expert knowledge for analysis, ensuring better alignment with the execution environment.
+Additionally, QUITE introduces a specialized reasoning agent, which forms the core of its multi-agent framework, incorporating a planning mechanism to sequence high-impact rewrite actions, maximizing performance improvements and preventing inefficiencies and regressions.
+Furthermore, QUITE incorporates a fine-grained hint injection technique into query rewrite process to further improve the query execution performance.
+To our knowledge, no other query rewrite systems have supported for these components.
 3. System Overview
-In this section, we present our system’s end-to-end pipeline (Section 3.1) and describe its core components (Section 3.2).
-3.1. System Pipeline
-The system pipeline is illustrated in Figure 3.
-❶ User provides the SQL query to be optimized along with the corresponding database configuration parameters (e.g., DBMS type, host address).
-❷ QUITE obtains metadata (e.g., database statistics, query information) to generate a reasoning chain.
-This chain, which contains rewrite proposals and corresponding rewritten SQL candidates, is then used to generate the rewritten SQL for downstream enhancement and validation.
-❸ QUITE consolidates the SQL candidates generated by the reasoning chains, constructs the prompt and selects and then applies global rewritten SQL enhancements.
-❹ QUITE validates the syntactical correctness and semantic equivalence of the rewritten SQL.
-❺ QUITE reassesses the rationality of the rewritten SQL to decide whether to proceed or to rewrite the SQL again referring to the rewrite history.
-In Query Hint Recommender, ❻ QUITE first reduces the hint search space by selecting and creating high-quality hints.
-Then, ❼ QUITE recommends and injects the most suitable query hints for the rewritten SQL.
-❽ Finally, QUITE outputs the final rewritten SQL together with optimization proposals.
-3.2. System Components
-QUITE has three components: Rewrite Middleware, LLM Agent-Based Query Rewriter, and Query Hint Recommender. They work as follows:
+We present the system architecture of QUITE  as shown in Figure 3. QUITE comprises three components: LLM Agent-Based Query Rewriter, Rewrite Middleware, and Query Hint Recommender.
 LLM Agent-Based Query Rewriter.
-The rewriter utilizes a training-free and feedback-aware LLM agent-based Finite State Machine to rewrite queries in multiple stages.
-First, we formalize query rewrite as a Markov Decision Process (MDP), treating each SQL form as a state, rewrite operations as actions, and cost‐reduction estimates as rewards to guide the LLM’s chain‐of‐thought toward lower‐cost rewrites (see Subsection 5.1).
-Second, we decompose the rewrite workflow into an FSM with three specialized agents, utilizing a multi‐agent collaboration strategy and constructing an FSM to ensure a stable workflow (see Subsection 5.2).
+The rewriter uses a training-free and feedback-aware LLM agent-based FSM to rewrite queries in multiple stages.
+First, ❶ a user provides the input SQL along with the corresponding database configuration (e.g., DBMS type, host address).
+Second, ❷ The rewriter obtains metadata (e.g., database statistics, query information) to generate a reasoning chain, which includes rewrite proposals and rewritten SQL candidates.
+This chain is then used to generate the rewritten SQL for further enhancement and validation.
+Next, ❸ we consolidate the SQL candidates to construct the LLM prompt, and apply global SQL enhancements to find the optimal rewritten SQL.
+Finally, ❹ we validate the syntactical correctness and semantic equivalence of the rewritten SQL, and ❺ reassess the rationality of the rewritten SQL to decide whether to proceed or rewrite the SQL again based on the rewrite history.
 Rewrite Middleware.
-The toolkit prepares three specialized tools to enhance the rewrite capability of LLM agents.
-First, we construct a structured knowledge base using a multi-source query rewrite knowledge integration pipeline to guide the LLM agent to generate high-quality rewrite proposals (see Section 5.1).
-Second, we develop a hybrid SQL corrector incorporating an LLM agent to validate and correct the syntactical correctness and semantic equivalence of the rewritten SQL (see Section 5.2).
-Third, we design an agent memory buffer that can extract key information from previous agent conversations, thus reducing agent communication overhead and LLM hallucination in the rewrite process (see Section 5.3).
+The toolkit provides three specialized tools to boost LLM rewrite ability.
+First, we construct a structured knowledge base using a multi-source integration to guide high-quality rewrite proposals (see Section 5.1).
+Second, a hybrid SQL corrector then validates the syntactical correctness and semantic equivalence of each candidate rewrite, reassessing rationality against rewrite history and triggering further refinement when necessary (see Section 5.2).
+Third, we design an agent memory buffer that captures key context from previous interactions to reduce communication overhead and LLM hallucination (see Section 5.3).
 Query Hint Recommender.
 The recommender explores the query hints in two steps.
-First, we carefully construct a promising query hint base, thereby reducing the search space (see Section 6.1).
-Second, we select appropriate hints for the rewritten query to improve its execution plan (see Section 6.2).
-Finally, the selected hints are injected into the rewritten SQL to produce the final output.
+❻ First, we carefully construct a promising query hint base, thereby reducing the hint search space and enabling faster, more accurate hint selection (see Section 6.1).
+❼ Second, we leverage LLM agent with data statistics to select appropriate fine-grained hints for the rewritten query to improve its execution performance (see Section 6.2).
+❽ Finally, the selected hints are injected into the rewritten SQL to produce the final output.
 4. LLM Agent-Based Query Rewriter
-In this section, we present our LLM agent–based query rewriter and describe how to construct an efficient, robust query‐rewriting workflow using LLM agents.
+In this section, we present our LLM agent-based query rewriter and describe how to construct an efficient, robust query‐rewriting workflow using LLM agents.
 Figure 4. Reasoning Chain of MDP-based Reasoning Agent
@@
 Reasoning LLM agents (e.g., DeepSeek-R1 (Guo et al., 2025)), capable of chain-of-thought planning and multi-step logical inference, appear well-suited to this task.
-However, our experiments show they still produce 13.0% nonequivalent outputs and yield minimal execution improvements (Section 7.2).
-We observe that query rewrite mirrors a Markov Decision Process, where states are query forms, actions are rewrite steps, and rewards are cost reductions (Puterman, 1990; Liu et al., 2023; Wang, 2025).
+However, our experiments show that they still produce 13.0% nonequivalent outputs and yield minimal execution improvements (Section 7.2).
+We observe that query rewrite mirrors a Markov Decision Process (MDP), where states are query forms, actions are rewrite steps, and rewards are cost reductions (Puterman, 1990; Liu et al., 2023; Wang, 2025).
 By framing rewrite in this way, we provide the reasoning LLM agent with a clear, stepwise optimization goal that guides each inference step.
@@
 Single-agent approaches often produce suboptimal or invalid rewrites because a single agent is responsible for managing the entire rewrite pipeline (e.g., propose rewrites, verify validity, and assess efficiency), which overloads the model’s context capacity and ultimately increases the risk of hallucination.
-Inspired by multi-agent frameworks that assign specialized agents to distinct tasks (Hong et al., 2023), we break down query rewrite into discrete stages, each handled by a dedicated LLM agent.
+Inspired by multi-agent frameworks that assign specialized agents to distinct tasks (Hong et al., 2023), we break down query rewrite into discrete stages, each is handled by a dedicated LLM agent.
 To ensure the multi-agent system’s stability and correctness (Han et al., 2024), we implement a Finite State Machine (FSM) to orchestrate stable stage transitions.
@@
 The Decision stage determines whether to continue or terminate the rewrite process.
+The Termination stage ends the rewrite process and outputs the rewritten SQL with rewrite proposals.
 4.2.2. Specialized LLM Agents
@@
 Since this candidate may introduce new structural complexities, it cannot be used directly for verification.
-Therefore, the Rewrite agent performs an SQL Enhancement Process—using the prompt template below—to identify further rewrite opportunities.
+Therefore, the rewrite agent undertakes an SQL enhancement process to uncover additional rewrite opportunities.
+The prompt template for this process is shown below:
 […]
@@
 If you think the SQL candidate is well-optimized, just return the original version as the rewritten SQL.
+This template provides a concise, user-curated checklist of the effective SQL enhancement transformations.
+Guided by this checklist, the rewrite agent revisits the selected SQL candidate to identify any remaining optimization opportunities.
+If none are found, it returns the original query unchanged.
 Assistant Agent. The Assistant agent uses the hybrid SQL corrector (Section 5.2) to ensure the syntactically correctness and semantically equivalence of queries. This is performed in two steps:
@@
 If the query remains uncorrected after these attempts, the current FSM iteration is aborted, and control returns to the reasoning stage.
-  • Equivalence correction. Once syntax‐correct, the agent submits the query pair $\langle Q_{\text{orig}},Q_{\text{new}}\rangle$ to verify the equivalence and correct nonequivalent queries.
+  • Equivalence correction.
+Once syntax error is corrected, the agent submits the query pair $\langle Q_{\text{orig}},Q_{\text{new}}\rangle$ to verify the equivalence and correct nonequivalent queries.
 This two-phase process ensures that only valid, semantically equivalent rewrites advance to the Decision stage.
@@
 If True, the Decision agent terminates the FSM process; If False, it retrieves higher-confidence rewrite proposals from the knowledge base, then stores the current query, report, and retrieved knowledge in an agent shared buffer for the next FSM iteration.
-We deliberately avoid introducing knowledge base retrieval in the initial state, because reasoning models have input content limitations—excessive guidance information can reduce the generation freedom, potentially compromising the quality of reasoning proposal chains (Guo et al., 2025).
+We deliberately avoid introducing knowledge base retrieval in the initial state, because reasoning models have input content limitations—excessive guidance information can reduce the generation freedom, potentially compromising the quality of reasoning proposal chain (Guo et al., 2025).
 Finally, the FSM enters the Termination stage and outputs the final rewritten query along with its comprehensive optimization report.
@@
 5.1. Structured Knowledge Base
-Preparing the knowledge base involves the following steps: (i) collecting SQL rewrite knowledge from various resources, (ii) filtering untrustworthy content, (iii) enhancing the knowledge, and (iv) retrieving relevant knowledge for specific tasks.
+Since DBAs encode numerous scenario-specific rewrite strategies that LLMs cannot learn from sparse examples and that fail to generalize across diverse query contexts, we consolidate these expert rules into a structured, scenario-indexed knowledge base.
+To this end, preparing the knowledge base involves the following steps: (i) collecting SQL rewrite knowledge from various resources, (ii) filtering untrustworthy content, (iii) enriching the entries, and (iv) retrieving the most relevant rules for each rewrite task.
 The Knowledge Base $\mathcal{K}$ is defined as follows:
@@
 These embeddings are then used to query a pre-indexed knowledge base of extracted points, retrieving the top three matches via cosine similarity.
-Finally, an LLM evaluates these matches to confirm they effectively explain the rewrite strategy underlying the Q&A unit.
+Finally, an LLM evaluates these matches to confirm that they effectively explain the rewrite strategy underlying the Q&A unit.
 Suitable matches are incorporated, enriching Q&A units with concise explanations that aid future retrieval.
 Step 4: Retrieving relevant knowledge.
-We classify these units into four categories: join optimization, constant folding, predicate simplification and other, which capture the essential aspects of query rewriting and provides a foundational structure for enabling precise retrieval of information.
+We classify these units into four categories: join optimization, constant folding, predicate simplification and others, which capture the essential aspects of query rewriting and provides a foundational structure for enabling precise retrieval of information.
 Since knowledge bases are often context-specific, we use the BM25 algorithm (Robertson et al., 2009) to retrieve the most relevant documents by computing the document relevance.
@@
 However, relying solely on LLMs carries inherent risks.
-Therefore, we propose a hybrid approach that integrates traditional verification tools with LLM capabilities to achieve broader and more reliable equivalence judgments in two stages.
+Therefore, we integrate traditional verification tools and LLM capabilities in a two-stage process, forming a hybrid approach that achieves broader and more reliable equivalence judgments.
   • Tool-based Verification.
@@
 Query Hint Grammar Description
-No Hash Join /*+ NoHashJoin(table table[ table...]) */ Prevents the use of hash joins for specified tables.
+Hash Join /*+ HashJoin(table table[ table...]) */ Forces the use of hash joins for specified tables.
 No Nested Loop Join /*+ NoNestLoop(table table[ table...]) */ Prevents the use of nested loop joins for specified tables.
@@
 Row Correction /*+ Rows(a b #10) */ Sets the number of rows for the join result.
-NO_MATERIALIZE /*+ NO_MATERIALIZE(table) */ Inlines the CTE to avoid materialization overhead.
+NOT MATERIALIZE /*+ NOT MATERIALIZE(table) */ Inlines the CTE to avoid materialization overhead.
 6. Query Hint Recommender
@@
 To exert more fine-grained control over plan selection, we propose embedding query hints (Marcus et al., 2021) within rewritten queries.
-In this work, we introduce a curated hint base and use LLMs to adaptively select and incorporate effective hints into rewritten queries.
+In this work, we introduce a dedicated hint base and use LLMs to adaptively select and incorporate effective hints into rewritten queries.
 6.1. Hint Base Construction
@@
 This section details our approach to using LLMs for identifying operations that could benefit from hints and then recommending appropriate ones.
-Unlike existing methods (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021) that recommend hints at the query level (based on global parameters like GUCs), where a single hint applies to all operations within that query, our strategy targets a more granular level.
+Unlike existing methods (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021) that recommend hints at the query level, where a single hint applies to all operations within that query, our strategy targets a more granular level.
 We found that query-level hints can be suboptimal, as a hint beneficial for one operation might negatively impact others.
 Therefore, we apply hints at a finer-grained level, such as to individual tables or specific operations within a query, to maximize positive impacts.
+Building on the validated rewrite of $Q^{\prime}$, we now turn to refining its physical execution plan through targeted hint injection.
 First, given the physical plan of $Q^{\prime}$, database semantics and data distribution statistics, we identify potential improvements that can be achieved through hints: (1) For every cardinality estimation in the plan, an LLM judges its reasonableness.
@@
 Testbed.
-Our experiments were conducted on a cloud server featuring a 32-core Intel Xeon Platinum 8352V CPU, 251 GB of RAM, a 942 GB SSD, and PostgreSQL v14.13.
+Our experiments were conducted on a server featuring a 32-core Intel Xeon Platinum 8352V CPU, 251 GB of RAM, a 942 GB SSD, and PostgreSQL v14.13.
 We utilize a NVIDIA A100-SXM4 GPU with 40 GB of memory to run the rewrite process of LLM-R2.
@@
 Baselines. We compare our system against the following state-of-the-art query rewrite baselines:
-  • LearnedRewrite (LR) (Zhou et al., 2021) uses a Monte Carlo Tree Search (MCTS) algorithm with a learned cost estimation model to explore a SQL policy tree of rewrite rule orders.
+  • LearnedRewrite (LR) (SIGMOD 2022)  (Zhou et al., 2021) uses a Monte Carlo Tree Search (MCTS) algorithm with a learned cost estimation model to explore a SQL policy tree of rewrite rule orders.
 We adopt the original implementation and use the cost estimation model provided in their official code repository.
-  • LLM-R2 (Li et al., 2024) is a rule-based query rewrite system built on Apache Calcite (Begoli et al., 2018).
+  • LLM-R2 (VLDB 2025) (Li et al., 2024) is a rule-based query rewrite system built on Apache Calcite (Begoli et al., 2018).
 It uses LLM’s In-Context Learning (ICL) ability to select query rewrite rules according to high quality demonstration queries.
+  • R-bot (Sun et al., 2024) is an LLM-based system that leverages an embedded retrieval-augmented generation (RAG) knowledge base and reflective reasoning to select promising rewrite rules.
   • LLM Agent rewrites queries using an individual LLM agent. To ensure a fair comparison, this LLM agent can access target database statistics and our SQL corrector for self-reflection.
-  • R-bot (Sun et al., 2024) is an LLM-based system that leverages an embedded retrieval-augmented generation (RAG) knowledge base and reflective reasoning to select promising rewrite rules.
 Table 3. Different Methods on Query Latency (s)
@@
 LLM-R2(DS-R1) 56.73 9.25 29.33 300.00 9.16 3.84 8.21 24.21 16.90 2.05 7.71 49.91
+R-Bot (GPT-4o) 32.64 9.48 26.25 172.20 21.75 4.26 8.80 119.19 22.21 2.34 8.35 122.14
+R-Bot (Claude-3.7) 33.89 9.71 29.32 195.51 19.89 4.27 8.99 112.24 23.38 2.60 8.64 122.95
+R-Bot (DS-R1) 33.70 9.88 28.38 176.90 21.21 4.44 9.11 113.91 22.71 2.52 8.43 122.89
 LLM Agent (GPT-4o) 69.03 10.00 27.02 300.00 33.44 4.39 9.65 300.00 15.47 2.03 7.32 35.34
@@
 LLM Agent (Claude-3.7) 57.46 11.04 36.99 300.00 25.32 4.38 8.84 300.00 15.67 1.59 7.68 44.54
-R-Bot (GPT-4o) 32.64 9.48 26.25 172.20 21.75 4.26 8.80 119.19 22.21 2.34 8.35 122.14
-R-Bot (Claude-3.7) 33.89 9.71 29.32 195.51 19.89 4.27 8.99 112.24 23.38 2.60 8.64 122.95
-R-Bot (DS-R1) 33.70 9.88 28.38 176.90 21.21 4.44 9.11 113.91 22.71 2.52 8.43 122.89
 […]
@@
 Notably, the improvement margin of QUITE over LR is substantially larger for the DSB dataset compared to the TPC-H and Calcite datasets.
-This disparity arises primarily because: (1) The LR we use is trained on the TPC-H, LR can proficiently identify effective rewrite rules and navigate their constrained search spaces.
+This is primarily because: (1) The LR we use is trained on the TPC-H, LR can proficiently identify effective rewrite rules and navigate their constrained search spaces.
 This existing proficiency limits the remaining optimization potential, thereby reducing the margin for QUITE to demonstrate substantial further improvements.
@@
 6.08 3.78 7.98 21.43
-LLM-Agent-based Query Rewrite FSM To evaluate the impact of the FSM framework within our system, we performed two modifications: (1) the loop responsible for equivalence checking was removed, and (2) the decision stage was eliminated.
+LLM-Agent-based Query Rewrite FSM.
+To evaluate the impact of the FSM framework within our system, we performed two modifications: (1) the loop responsible for equivalence checking was removed, and (2) the decision stage was eliminated.
 These changes transformed the process into a linear, end-to-end pipeline while retaining all components and agent tools involved in query rewriting.
@@
 QUITE (Rewrite on SF=30) 5.72 0.50 1.17 5.17 30.50 9.96 24.29 96.68 40.65 16.89 37.96 131.97
-Query Hint Injection We evaluate the impact of hint injection on queries rewritten by QUITE.
+Query Hint Injection.
+We evaluate the impact of hint injection on queries rewritten by QUITE.
 As shown in Table 3, the insertion of hints leads to a further decrease in average execution times by 1.8%, 3.8%, and 0.3% on TPC-H, DSB, and Calcite, respectively.

### v3

@@
-QUITE: A Query Rewrite System Beyond Rules with LLM Agents
+QUITE: A Query Rewrite System Beyond Rules via LLM Agents
 Abstract.
@@
 1. Introduction
-Inefficient SQL queries remain a long-standing challenge in many database applications, often arising from poorly written queries by inexperienced users or automatically generated systems (Li, 2019; Zhou et al., 2020; Pirahesh et al., 1992).
-Query rewrite is a promising approach that transforms an original SQL into a semantically equivalent form with better performance (Finance and Gardarin, 1991a; Pirahesh et al., 1992; Wang et al., 2022).
-An effective rewritten query can improve the execution time of a poorly written SQL query by several orders of magnitude (Wang et al., 2022; Zhou et al., 2021).
-Existing query rewrite approaches typically rely on predefined rewrite rules (Begoli et al., 2018; Graefe and McKenna, 1993; PostgreSQL Global Development Group, 2024; Zhou et al., 2021, 2023; Bai et al., 2023; Li et al., 2024; Sun et al., 2024).
-Heuristic-based methods apply these rules in a fixed order derived from practical experience (Begoli et al., 2018; Graefe and McKenna, 1993; PostgreSQL Global Development Group, 2024).
+Inefficient SQL queries remain a long-standing challenge in many database applications, often arising from poorly written queries by inexperienced users or automatically generated systems (cloud; database; extensible).
+Query rewrite is a promising approach that transforms an original SQL into a semantically equivalent form with better performance (query-rewrite; extensible; wetune).
+An effective rewritten query can dramatically improve the execution time of a poorly written SQL query by several orders of magnitude (wetune; lr).
+Existing query rewrite approaches typically rely on predefined rewrite rules (apache; volcano; pg; lr; lr_demo; querybooster; llm-r2; r-bot).
+Heuristic-based methods apply these rules in a fixed order derived from practical experience (apache; volcano; pg).
 However, such fixed orders often fail to generalize across different environments and query patterns.
-Thus, machine learning (ML)-based methods train models on rewrite histories to predict which rules to use and the order in which to apply them (Zhou et al., 2021, 2023).
-Recently, some approaches have explored the use of Large Language Models (LLMs) to select rewrite rules and determine their application order based on domain knowledge (Li et al., 2024; Sun et al., 2024).
+Thus, machine learning (ML)-based methods train models on rewrite histories to predict which rules to use and the order in which to apply them (lr; lr_demo).
+Recently, some approaches have explored the use of Large Language Models (LLMs) to select rewrite rules and determine their application order based on domain knowledge (llm-r2; r-bot).
 Figure 1. An Example of Using Strategies Beyond Rewrite Rules (Example 1.1) and Using Context-Aware Analysis (Example 1.2)
 Figure 2. An Example of Using Query Intent (Example 1.3)
-However, the above approaches can handle only a small subset of queries and may lead to performance regressions (Zhou et al., 2021; Li et al., 2024; Sun et al., 2024).
+However, the above approaches can handle only a small subset of queries and may lead to performance regressions (lr; llm-r2; r-bot).
 This limitation stems from the intrinsic challenges of rule-based query rewrite.
-First, the rich features of SQL and the nuances of semantics make it rather difficult to discover and verify new rewrite rules (Pirahesh et al., 1992; Chu et al., 2017b; Wang et al., 2022).
-Even state-of-the-art rule discovery system (Wang et al., 2022) can only support a narrow subset of operators.
-Second, fixed rewrite rules rely on pattern matching and therefore are fundamentally unable to optimize unseen or complex query patterns (Leis et al., 2015; Bai et al., 2023; Dong et al., 2023).
-Third, many effective rewrite strategies cannot be expressed by rewrite rules (Dong et al., 2023).
+First, the rich features of SQL and the nuances of semantics make it rather difficult to discover and verify new rewrite rules (extensible; hottsql; wetune).
+Even state-of-the-art rule discovery system (wetune) can only support a narrow subset of operators.
+Second, fixed rewrite rules rely on pattern matching and therefore are fundamentally unable to optimize unseen or complex query patterns (Leis2015; querybooster; slabcity).
+Third, many effective rewrite strategies cannot be expressed by rewrite rules (slabcity).
 For example, the Common Table Expressions (CTE) conversion in Example 1.1 cannot be captured by rewrite rules.
-This is because such conversion requires explicit control over evaluation order and sub-plan sharing, which is beyond the scope of relational-algebra rewrite rules (Graefe, 1993; Wang et al., 2022; Begoli et al., 2018).
+This is because such conversion requires explicit control over evaluation order and sub-plan sharing, which is beyond the scope of relational-algebra rewrite rules (graefe1993query; wetune; apache).
 Given these limitations, we revisit the query rewrite problem from the perspective of human experts and identify three key capabilities that humans exhibit, but rule-based systems lack.
-(1) Humans can perform rewrites that follow rules but go beyond the expressiveness of pattern matching (Example 1.1).
+(1) Humans can perform rewrites that follow certain rules but exceed the expressiveness of standard pattern matching, and in some cases, go beyond what can be captured by formal rewrite rules (Example 1.1).
 (2) Humans can conduct context-aware rewrites by considering the underlying data distributions and cost estimates provided by the query optimizer (Example 1.2).
@@
 Query Rewrite Beyond Rewrite Rules.
-Figure 1 (a) is a simplified query from the DSB benchmark (Ding et al., 2021).
+Figure 1 (a) is a simplified query from the DSB benchmark (ding2021dsb).
 To avoid repeated calculation of nested subqueries, Figure 1 (b) rewrites them using Common Table Expressions (CTEs).
@@
 Context-Aware Query Rewrite.
-In Figure 1 (c), instead of directly joining two large tables (the sale with 10 M rows and the ttime with 1.4 M rows), we first joins the sale table with the prod table (100 K rows).
+In Figure 1 (c), instead of directly joining two large tables (the sale with 10 M rows and the ttime with 1.4 M rows), we first join the sale table with the prod table (100 K rows).
 This highly selective join would produce an intermediate result of only  100 K rows.
@@
 Query Rewrite based on Query Intent.
-Figure 2 (a) shows a real-world query from Apache Calcite (apa, 2021).
+Figure 2 (a) shows a real-world query from Apache Calcite (apache_calcite_test).
 Rule-based rewrite cannot optimize this query (Figure 2 (b)).
-Identifying that the query intent is to sum salaries directly from the EMP table, human can simplify the original query by eliminating unnecessary joins and sub-queries that do not contribute to the intent of the query, achieving a 2.17 $\times$ execution improvement in our experimental setup.
-Although human experts can produce significantly better SQL rewrites than rule-based methods, manual rewrite suffers from poor scalability, especially when there are thousands or even more poorly written queries to optimize in cloud environments.
+Identifying that the query intent is to sum salaries directly from the EMP table, humans can simplify the original query by eliminating unnecessary joins and sub-queries that do not contribute to the intent of the query, achieving a 2.17 $\times$ execution improvement in our experimental setup.
+Although human experts can produce significantly better SQL rewrites than rule-based methods, manual rewriting does not scale well, especially when there are millions of poorly written queries to be optimized in cloud environments.
 Recently, LLMs have demonstrated nearly human-level semantic and reasoning abilities, making them a natural choice for query rewrite.
@@
 Ensuring Optimized Rewrites.
-Determining whether a rewritten query improves performance depends on multiple factors, including query intent, data distribution, optimizer estimates, system states, index configurations, and others.
-These contextual factors are typically unavailable to LLMs, and enabling LLMs to effectively utilize such information remains a non-trivial problem (Lao et al., 2024a; Bai et al., 2024; Tan et al., 2025).
+Determining whether a rewritten query improves performance depends on multiple factors, including query intent, data distribution, optimizer estimates, system states, and others.
+These contextual factors are typically unavailable to LLMs, and enabling LLMs to effectively utilize such information remains a non-trivial problem (gptuner; bai-etal-2024-longbench; llmqo).
 To address the limitations of rule-based query rewrite, the poor scalability of manual rewrite, and the unreliable nature of LLMs, we propose QUITE, a training-free and feedback-aware system that smartly leverages LLMs to rewrite SQL queries into semantically equivalent forms with significantly improved performance.
 QUITE supports a broader range of query patterns and rewrite strategies, while maintaining good scalability.
+It is primarily designed for long-running OLAP workloads, where execution time dominates and the cost of query rewriting is negligible.
 First, to enable LLMs to capture query data characteristics and leverage database feedback, we design a multi-agent framework controlled by a finite state machine (FSM) to iteratively refine the rewrite process.
 We decompose the complex rewrite process into a set of subtasks, each handled by a specialized LLM agent to reduce hallucinations and improve rewrite ability (addressing C1, C2).
-Second, to enhance the ability of LLMs to generate high-quality query rewrites, we provide a rewrite middleware that includes: (1) a structured knowledge base extracted from official documentation and query rewrite discussions on web forums, (2) a hybrid SQL corrector to ensure query equivalence, and (3) an agent memory buffer that manages important rewrite contexts (addressing C1, C2).
-Third, we employ a novel hint injection technique to generate better execution plans for rewritten queries.
+Second, to enhance the ability of LLMs to generate high-quality query rewrites, we provide a rewrite middleware that includes: (1) a structured knowledge base extracted from official documentation and query rewrite discussions on web forums, (2) a hybrid SQL corrector that combines tool-based and LLM-based equivalence checking to ensure query equivalence, Third, we employ a novel hint injection technique to generate better execution plans for rewritten queries.
 This ensures that query execution exactly follows the optimized plan recommended by LLM agents, avoiding unintended modifications by the query optimizer due to inaccurate cost estimations.
 Additionally, this technique enables further improvements by incorporating physical-level optimizations (addressing C2).
-We extensively evaluate QUITE against state-of-the-art query rewrite systems on widely used benchmarks (e.g., TPC-H, DSB, and Calcite) and show that our system delivers up to a 21.9% speedup in query execution performance over the best-performing alternative.
+We extensively evaluate QUITE against state-of-the-art query rewrite systems on widely used benchmarks (e.g., TPC-H, DSB, and Calcite), which shows that our system achieves up to a 21.9% speedup in query execution performance over the best-performing alternative.
 Moreover, it delivers 24.1% additional rewrites compared to that alternative, covering query cases that previous systems could not handle.
@@
   • We propose QUITE, a training-free and feedback-aware query rewrite system that leverages LLM agents to support a broader range of query patterns and rewrite strategies than existing methods (Section 3).
-  • We design a multi-agent framework controlled by a finite state machine to enable LLMs to use external tools and enhance the rewrite process with database feedback (Section 4).
+  • We design a multi-agent framework controlled by a finite state machine to enable LLM agents to use external tools and enhance the rewrite process with database feedback (Section 5).
   • We develop a rewrite middleware including a structured knowledge base, a hybrid SQL corrector, and an agent memory buffer to enhance the rewrite capability of LLM agents (Section 4).
@@
 2. Preliminaries and Related Work
+In this section, we first formally define the query rewriting problem and then discuss related work on query hints and LLMs.
 Figure 3. System Overview of QUITE
@@
 (1) Discovery of new rules.
-Approaches such as WeTune (Wang et al., 2022) automatically generate and verify logical plan transformations, but they are limited to specific operator types and algebraic equivalences.
-Equality saturation (Nandi et al., 2021) and user-guided tools like QueryBooster (Bai et al., 2023) lack full autonomy and broad applicability.
-Recently, GenRewrite (Liu and Mozafari, 2024) uses LLMs to suggest rules in natural language.
-However, it still struggles with optimal rewrite selection because executing all candidates to measure latency incurs prohibitive overhead and its off-the-shelf verifiers permit equivalence checking for only a subset of rewrite pairs (Liu and Mozafari, 2024; Dong et al., 2023).
+WeTune (wetune) automatically generates and verifies logical plan transformations, but it handles only a limited set of operator types and algebraic rules.
+Equality saturation techniques (10.1145/3485496) and user-guided systems like QueryBooster (querybooster) still require manual intervention and cover only narrow classes of queries.
+More recently, GenRewrite (genrewrite) leverages LLMs to describe rewrite rules in natural language (NLR2s), enabling more flexible rule specification.
+In practice, its effectiveness depends on the coverage of available equivalence checks, and validating a large number of candidate rewrites through execution can introduce substantial overhead (genrewrite; slabcity).
 (2) Effective use of existing rules.
-A rule-based rewrite framework must answer two key questions: which rules to include in the rule set and how to select among applicable rules during rewriting.
-LearnedRewrite (Zhou et al., 2021) uses Monte Carlo Tree Search (MCTS) with learned cost models to explore the space of rule applications, but its performance depends on the accuracy of cost model estimation.
-LLM-R2 (Li et al., 2024) uses a trained selection model to retrieve the past rewrite query and rule demonstrations similar to the input SQL, and leverages them as an in-context prompt for the LLM to select promising rewrite rules.
-However, it still incurs high costs by executing every query demonstration on each rule set and faces distribution shifts between the demonstration and input SQL, since structurally similar queries may differ in row count, leading to suboptimal rule selection.
+LearnedRewrite (lr; lr_demo) uses Monte Carlo Tree Search (MCTS) with learned cost models to explore the space of rule applications, but its performance depends on the accuracy of cost model estimation.
+LLM-R2 (llm-r2) retrieves similar past rewrites using a pre-trained selector and employs them as in-context examples for rule selection.
+R-Bot (r-bot) adopts retrieval-augmented generation to recommend rewrite rules through evidence retrieval and step-by-step reasoning.
+These approaches largely retrieve and apply rewrite knowledge based on query structure and then select from predefined rule sets.
+Consequently, they may not fully capture dataset-dependent factors such as database scale or column statistics, and their rule-centric formulation can miss effective expert strategies that are hard to express as explicit rewrite rules.
 2.2. Query Hints
-Query hints are SQL extensions that provide instructions to the database’s query engine to influence the selection of execution plans (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021).
+Query hints are SQL extensions that provide instructions to the database’s query engine to influence the selection of execution plans (bao; autosteer; fastgres; steeringqo).
 These hints allow users to exert precise control over how the database optimizer chooses physical operators and access methods.
-As shown below for pg_hint_plan extension (NTT OSS Center DBMS Development and Support Team, 2012), the hint /*+ HashJoin(employees departments) */ forces the query engine to use a hash join for the $employees\bowtie departments$ operation, while the /*+ SeqScan(employees)*/ enforces a full table scan of the employees relation.
-These hints can override the optimizer’s default cost-based plan selection.
-HashJoin(employees departments)
-SeqScan(employees)
+As shown below for pg_hint_plan extension (pg_hint_plan), the hint /*+ HashJoin(employees departments) */ forces the query engine to use a hash join for the $employees\bowtie departments$ operation.
+This hint can override the optimizer’s default cost-based plan selection.
+/*+ HashJoin(employees departments) */
 SELECT *
@@
 WHERE e.salary > 50000;
-Although existing studies on query hints (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023) have shown promising results, they typically focus on boolean hints applied at the query level (e.g., the enable_nestloop configuration parameter in PostgreSQL), which apply globally and can affect subsequent workload queries.
-In contrast, pg_hint_plan injects fine‐grained hints at the level of individual joins or cardinality estimates, allowing the use of multiple hint types (e.g., cardinality adjustments) at a more fine-grained level (e.g., specifying join algorithms for particular table pairs) while requiring a more precise selection mechanism.
+Prior works on query hint selection (bao; autosteer; fastgres) apply boolean hints to entire queries, which often leads to suboptimal performance because different operators within a query may require different hints.
+Proto-X (zhang2024holon) improves upon this by assigning hints at the operator level.
+However, it operates in an offline setting that requires exploring multiple configurations and repeatedly executing the complete workload, whereas query rewriting does not require repeated workload execution.
+UniTune (zhang2023unified) supports query rewriting through LearnedRewrite (lr) but does not incorporate hints.
+In contrast, QUITE employs pg_hint_plan to enable fine-grained, operator-level control (e.g., selection of join algorithms and adjustment of cardinalities) and jointly optimizes both query rewriting and hint selection.
 2.3. LLM and LLM Agents
 LLM for Database Tasks.
-LLMs are increasingly applied to database tasks like text-to-SQL (Gao et al., 2023), SQL workload generation (Lao and Trummer, 2025), data processing (Albalak et al., 2024), and database optimization (Lao et al., 2024a; Zhao et al., 2025; Lao et al., 2024b; Tan et al., 2025; Lao et al., 2025), with recent work focusing on query optimization (Li et al., 2024; Sun et al., 2024; Liu and Mozafari, 2024).
+LLMs are increasingly applied to database tasks text-to-SQL (text2sql2; codes; omnisql; xiyansql), SQL workload generation (SQLBarber; lao2025sqlbarberleveraginglargelanguage), and database optimization (gptuner; llmidx; gptuner-demo; llmqo; gptuner-record), with recent work focusing on query optimization (llm-r2; r-bot; genrewrite).
 Although these methods show that LLMs can understand and modify SQL, the inherent complexity of query rewrite presents substantial challenges.
 LLM Agents.
-The primary advantage of LLM agents over LLMs is their ability to collect feedback from the execution environment (e.g., database execution metrics) and dadapt their subsequent reasoning and outputs accordingly (Wu et al., 2025).
-They have proven particularly effective in complex problem solving—for example, SWE‐agent in software engineering (Yang et al., 2024) and ChemCrow in chemistry (Bran et al., 2023).
-QUITE applies this LLM agents methodology to query rewrite by gathering DBMS feedback and adjusting each query candidate accordingly.
-This performs a feedback-driven refinement, enabling reliable and advanced query rewrite that extends beyond predefined rules or the capabilities of standalone LLMs.
-2.4. Comparison with Other Works
-Instead of focusing on finding optimal rewrite rules with limited coverage (Zhou et al., 2021; Li et al., 2024; Sun et al., 2024), QUITE addresses the core aspects of query rewriting by identifying semantic and physical rewrite bottlenecks through real-time interaction between LLM agents and the database, enhancing the detection of performance issues in the input SQL.
-Furthermore, QUITE generates a reasoning chain with natural language rewrite proposals and rewrites the SQL through multi-agent collaboration within a carefully designed Finite State Machine (FSM) pipeline and agent tools, ensuring equivalence and effectiveness.
-This approach enhances adaptability and efficiency in context-aware rewrites, compared to rule-based systems that miss complex query patterns and context-specific optimizations.
-In contrast to GenRewrite (Liu and Mozafari, 2024), which directly utilizes LLMs to generate rewrite rules for rewriting, QUITE considers key factors such as the SQL’s physical plan, the database distribution, and the domain expert knowledge for analysis, ensuring better alignment with the execution environment.
-Additionally, QUITE introduces a specialized reasoning agent, which forms the core of its multi-agent framework, incorporating a planning mechanism to sequence high-impact rewrite actions, maximizing performance improvements and preventing inefficiencies and regressions.
-Furthermore, QUITE incorporates a fine-grained hint injection technique into query rewrite process to further improve the query execution performance.
-To our knowledge, no other query rewrite systems have supported for these components.
+The primary advantage of LLM agents over LLMs is their ability to collect feedback from the execution environment (e.g., database execution metrics) and adapt their subsequent reasoning and outputs accordingly (wu2025agenticreasoningreasoningllms).
+They have proven particularly effective in code generation (liu2023your; yang2024sweagentagentcomputerinterfacesenable), automated diagnostics (d-bot; mdagent; thompson2023large), and knowledge management (d-bot; asai2024self).
+Moreover, collaborative deployments of multiple LLM agents often outperform single‑agent systems on complex reasoning and planning tasks (multi1; multi2; multi3; multi4).
+QUITE applies this multi-agent methodology to query rewrite by gathering DBMS feedback and adjusting each query candidate accordingly.
 3. System Overview
-We present the system architecture of QUITE  as shown in Figure 3. QUITE comprises three components: LLM Agent-Based Query Rewriter, Rewrite Middleware, and Query Hint Recommender.
+We present the architecture of QUITE  as shown in Figure 3. It comprises three components: LLM Agent-Based Query Rewriter, Rewrite Middleware, and Query Hint Recommender.
 LLM Agent-Based Query Rewriter.
 The rewriter uses a training-free and feedback-aware LLM agent-based FSM to rewrite queries in multiple stages.
-First, ❶ a user provides the input SQL along with the corresponding database configuration (e.g., DBMS type, host address).
+First, ❶ the user provides the input SQL along with the corresponding database configuration (e.g., DBMS type, host address).
 Second, ❷ the rewriter obtains metadata (e.g., database statistics, query information) to generate a reasoning chain, which includes rewrite proposals and rewritten SQL candidates.
@@
 Next, ❸ we consolidate the SQL candidates to construct the LLM prompt, and apply global SQL enhancements to find the optimal rewritten SQL.
-Finally, ❹ we validate the syntactical correctness and semantic equivalence of the rewritten SQL, and ❺ reassess the rationality of the rewritten SQL to decide whether to proceed or rewrite the SQL again based on the rewrite history.
+Finally, ❹ we validate the syntax correctness and semantic equivalence of the rewritten SQL, and ❺ verify the correctness and effectiveness of the rewritten query to decide whether to proceed to Step ❻ or further rewrite this query based on the rewrite history.
 Rewrite Middleware.
 The toolkit provides three specialized tools to boost LLM rewrite ability.
-First, we construct a structured knowledge base using a multi-source integration to guide high-quality rewrite proposals (see Section 5.1).
-Second, a hybrid SQL corrector then validates the syntactical correctness and semantic equivalence of each candidate rewrite, reassessing rationality against rewrite history and triggering further refinement when necessary (see Section 5.2).
-Third, we design an agent memory buffer that captures key context from previous interactions to reduce communication overhead and LLM hallucination (see Section 5.3).
+First, we construct a structured knowledge base using a multi-source integration to guide high-quality rewrite proposals (Section 5.1).
+Second, a hybrid SQL corrector then validates the syntactical correctness and semantic equivalence of each candidate rewrite, using rewrite history to trigger further refinement when necessary.
+(Section 5.2).
+Third, we design an agent memory buffer that captures key context from previous interactions to reduce communication overhead and LLM hallucination (Section 5.3).
 Query Hint Recommender.
 The recommender explores the query hints in two steps.
-❻ First, we carefully construct a promising query hint base, thereby reducing the hint search space and enabling faster, more accurate hint selection (see Section 6.1).
-❼ Second, we leverage LLM agent with data statistics to select appropriate fine-grained hints for the rewritten query to improve its execution performance (see Section 6.2).
+❻ First, we construct a promising query hint base, thereby reducing the hint search space and enabling faster, more accurate hint selection (Section 6.1).
+❼ Second, we leverage an LLM agent with data statistics to select appropriate fine-grained hints for the rewritten query to improve its execution performance (Section 6.2).
 ❽ Finally, the selected hints are injected into the rewritten SQL to produce the final output.
+Figure 4. Reasoning Chain of MDP-based Reasoning Agent
 4. LLM Agent-Based Query Rewriter
 In this section, we present our LLM agent-based query rewriter and describe how to construct an efficient, robust query‐rewriting workflow using LLM agents.
-Figure 4. Reasoning Chain of MDP-based Reasoning Agent
 4.1. MDP-based Reasoning Agent
-Query rewrite is fundamentally a sequential decision problem (Finance and Gardarin, 1991b; Zhou et al., 2021), each transformation step changes the query structure and limits future optimization options.
-Reasoning LLM agents (e.g., DeepSeek-R1 (Guo et al., 2025)), capable of chain-of-thought planning and multi-step logical inference, appear well-suited to this task.
+Query rewrite is fundamentally a sequential decision problem (131472; lr), each transformation step changes the query structure and limits future optimization options.
+Reasoning LLM agents (e.g., DeepSeek-R1 (DeepSeek-r1)), capable of chain-of-thought planning and multi-step logical inference, appear well-suited to this task.
 However, our experiments show that they still produce 13.0% nonequivalent outputs and yield minimal execution improvements (Section 7.2).
-We observe that query rewrite mirrors a Markov Decision Process (MDP), where states are query forms, actions are rewrite steps, and rewards are cost reductions (Puterman, 1990; Liu et al., 2023; Wang, 2025).
+We observe that query rewrite mirrors a Markov Decision Process (MDP), where states are query forms, actions are rewrite steps, and rewards are cost reductions (mdp; liu2023reason; tutorial).
 By framing rewrite in this way, we provide the reasoning LLM agent with a clear, stepwise optimization goal that guides each inference step.
@@
 We model the query rewrite process as a Markov Decision Process $(\mathcal{S},\mathcal{A},\mathcal{T},r,\gamma)$.
-Starting from the original query $Q$, each state $S_{t}\in\mathcal{S}$ represents the current form of the query with any semantical-level rewrite attempts.
+Starting from the original query $Q$, each state $S_{t}\in\mathcal{S}$ represents the current form of the query with any semantic-level rewrite attempts.
 At each timestep $t$, the agent chooses either a refinement action $R_{t}$ (e.g., join reordering, predicate pushdown) or a terminal action $\mathcal{A}$ that emits the final rewritten query.
@@
 $S_{t+1}=\begin{cases}\mathcal{T}(S_{t},R_{t})=S_{t}+R_{t},&\text{if a refinement is applied},\\ \mathcal{T}(S_{t},a)=S_{t}+\mathcal{A},&\text{if rewriting terminates}.\end{cases}$
-To guide the search toward efficient rewritten queries, we define a domain‐specific reward function once $S_{t}$ transitions to $S_{t+1}$ under $R_{t}$:
+To guide the search toward efficient rewritten queries, we define a domain‐specific reward function once $S_{t}$ transitions to $S_{t+1}$ under a refinement action $R_{t}$:
 $r(S_{t},R_{t})\;=\;\mathrm{Cost}(S_{t})\;-\;\mathrm{Cost}(S_{t+1}),$
@@
 In practice, we realize the policy $\pi_{\theta}(a\mid S_{t})$ as a reasoning‐specialized LLM (e.g., DeepSeek-R1) that generates each action conditioned on the current query state.
-As illustrated in Figure 4, the agent transitions from the initial state $S_{0}$ through a sequence of refinement actions $\{R_{0},R_{1},\dots\}$, and finally applies the terminal action $A$ to emit the final rewritten query.
-Concretely, at each timestep $t$, the agent is prompted with the current SQL form $S_{t}$ and analyzes the query structure via its internal chain‐of‐thought, scores candidate refinements according to the expected cost reduction, and samples the highest‐scoring $R_{t}$.
+As shown in Figure 4, the agent transitions from the initial state $S_{0}$ through a sequence of refinement actions $\{R_{0},R_{1},\dots\}$, and finally applies the terminal action $A$ to emit the final rewritten query.
+Concretely, at each timestep $t$, the agent is prompted with the current SQL form $S_{t}$ and analyzes the query structure via its internal chain‐of‐thought, scores candidate refinements based on the expected cost reduction, and samples the highest‐scoring $R_{t}$.
 4.2. FSM-based Query Rewrite
 Single-agent approaches often produce suboptimal or invalid rewrites because a single agent is responsible for managing the entire rewrite pipeline (e.g., propose rewrites, verify validity, and assess efficiency), which overloads the model’s context capacity and ultimately increases the risk of hallucination.
-Inspired by multi-agent frameworks that assign specialized agents to distinct tasks (Hong et al., 2023), we break down query rewrite into discrete stages, each is handled by a dedicated LLM agent.
-To ensure the multi-agent system’s stability and correctness (Han et al., 2024), we implement a Finite State Machine (FSM) to orchestrate stable stage transitions.
+Inspired by multi-agent frameworks that assign specialized agents to distinct tasks (metagpt), we break down query rewrite into discrete stages, each of which is handled by a dedicated LLM agent.
+To ensure the multi-agent system’s stability and correctness (multi-agent-chanllenges), we implement a Finite State Machine (FSM) to orchestrate stable stage transitions.
 4.2.1. Query Rewrite Process Decomposition
@@
   • $\mathcal{G}=\{A_{1},...,A_{n}\}$ is a set of cooperative LLM-based agents, each with specialized capabilities $C_{i}$ for tasks.
-  • $\mathcal{M}:\mathcal{Q}\times\mathcal{E}\rightarrow\mathcal{S}$ is a transmission that breaks down the input query $Q\in\mathcal{Q}$ into intermediate states $s\in\mathcal{S}$.
+  • $\mathcal{M}:\mathcal{Q}\times\mathcal{E}\rightarrow\mathcal{S}$ is a transition that breaks down the input query $Q\in\mathcal{Q}$ into intermediate states $s\in\mathcal{S}$.
   • $\mathcal{F}:\mathcal{S}\times\mathcal{H}\rightarrow\mathcal{Q}^{\prime}$ is a transformation function that produces rewritten queries $Q^{\prime}\in\mathcal{Q}^{\prime}$ based on the current state and historical feedback $h\in\mathcal{H}$ from $\mathcal{E}$.
   • $\mathcal{R}:\mathcal{Q}^{\prime}\times\mathcal{E}\rightarrow\mathbb{R}^{+}$ is a reward function that evaluates rewrite quality through execution metrics and guides iterative refinement.
-Figure 5. The Workflow of FSM-based Query Rewrite
-Table 1. Specialized Agents in the Finite State Machine
-State Agent Agent Role
-$S_{1}$ MDP-based Reasoning Agent Generate Reasoning Chain
-$S_{1}$ Rewrite Agent Extract & Refine SQL candidates
-$S_{2}$ Assistant Agent Validate Syntax & Equivalence
-$S_{3}$ Decision Agent Score & Control Termination
 Specifically, the Agent takes action $O_{i}$ in state $s_{i}$, the environment provides feedback $O_{i}^{\prime}$, and the Agent updates its state and decides the next action based on the feedback as:
@@
 4.2.2. Specialized LLM Agents
-As shown in Table 1, we deploy three specialized LLM agents alongside the core MDP-based Reasoning Agent in Section 4.1 to drive the rewrite pipeline for defined FSM stages.
+As shown in Table 1, we deploy three specialized LLM agents alongside the core MDP-based Reasoning Agent in Section 4.1 to drive the rewrite pipeline for the defined FSM stages.
 Then we propose an LLM agent-based query rewrite methodology detailed in Algorithm 1:
@@
   • Verification stage ($S_{2}$): In this stage, the Assistant Agent validates syntactical correctness and semantic equivalence by using the hybrid SQL corrector (Lines 8-12).
-  • Decision stage ($S_{3}$): In this stage, the Decision Agent reassesses the rewritten SQL’s efficiency with comprehensive report, and determines whether the rewrite process should continue or terminate (Lines 13-19).
+  • Decision stage ($S_{3}$): In this stage, the Decision Agent reassesses the rewritten SQL’s efficiency with a comprehensive report and determines whether the rewrite process should continue or terminate (Lines 13-19).
+Figure 5. The Workflow of FSM-based Query Rewrite
+Table 1. Specialized Agents in the Finite State Machine
+State Agent Agent Role
+$S_{1}$ MDP-based Reasoning Agent Generate Reasoning Chain
+$S_{1}$ Rewrite Agent Extract & Refine SQL candidates
+$S_{2}$ Assistant Agent Validate Syntax & Equivalence
+$S_{3}$ Decision Agent Score & Control Termination
 1
@@
 9    $\langle Q^{E},\text{EarlyStop}\rangle\leftarrow\text{Enhance-Query}(\mathcal{M},Q^{P},\mathcal{A})$;
-10             $S\leftarrow\text{EarlyStop}\vspace{0.25pt}?\vspace{0.25pt}\textsc{Termination}:\textsc{Verification}$;
+10    $S\leftarrow\text{EarlyStop}\vskip 0.25pt?\vskip 0.25pt\textsc{Termination}:\textsc{Verification}$;
 11
@@
 20    $\mathcal{R}\leftarrow\text{Generate-Report}(Q^{E},Q_{0},\mathcal{M},T)$;
-[…]
+21    if $\text{Cost-Check}(Q^{E},\mathcal{R},\mathcal{M})$ then
 22       $Q^{*}\leftarrow Q^{E}$; $S\leftarrow\textsc{Termination}$;
@@
 32return $\langle Q^{*},\text{cost}(Q^{*}),\mathcal{A}\rangle$;
-Algorithm 1 LLM-Agent-based Query Rewrite Algorithm
+33
+Algorithm 1 LLM-based Query Rewrite Algorithm
 Rewrite Agent.
@@
 Therefore, the rewrite agent undertakes an SQL enhancement process to uncover additional rewrite opportunities.
-The prompt template for this process is shown below:
-[…]
-1. Push down predicates or use a CTE structure
-2. Simplify calculations and remove redundant predicates
-3. Apply other efficiency-preserving transformations
-If you think the SQL candidate is well-optimized, just return the original version as the rewritten SQL.
 This template provides a concise, user-curated checklist of the effective SQL enhancement transformations.
@@
 If none are found, it returns the original query unchanged.
-Assistant Agent. The Assistant agent uses the hybrid SQL corrector (Section 5.2) to ensure the syntactically correctness and semantically equivalence of queries. This is performed in two steps:
-  • Syntax correction.
-If syntax errors occur, it attempts up to $K_{\max}$ repairs.
-If the query remains uncorrected after these attempts, the current FSM iteration is aborted, and control returns to the reasoning stage.
+Assistant Agent. The assistant agent uses the hybrid SQL corrector (Section 5.2) to ensure the syntactic correctness and semantically equivalence of queries. This is performed in two steps:
+  • Syntax correction. If syntax errors occur, it attempts up to $K_{\max}$ repairs. If correction still fails, the current FSM iteration is aborted and control returns to the reasoning stage.
   • Equivalence correction.
-Once syntax error is corrected, the agent submits the query pair $\langle Q_{\text{orig}},Q_{\text{new}}\rangle$ to verify the equivalence and correct nonequivalent queries.
+Once the syntax error is corrected, the agent submits the query pair $\langle Q_{\text{orig}},Q_{\text{new}}\rangle$ to verify the equivalence and correct nonequivalent queries.
 This two-phase process ensures that only valid, semantically equivalent rewrites advance to the Decision stage.
-Although equivalence verification may return the original format, it avoids the far greater cost of returning nonequivalent SQL while preventing performance degradation from excessive loop iterations.
+Although equivalence verification may return the original input SQL, it avoids the far greater cost of returning nonequivalent SQL while preventing performance degradation from excessive loop iterations.
 Decision Agent. The decision agent implements a cost-aware decision mechanism to determine whether the rewrite process should terminate within budget constraints. It contains two steps:
@@
 The decision agent analyzes the rewritten query across four dimensions, including cost changes, plan characteristics, resource utilization, and other improvements.
-This multi-dimensional approach overcomes the limitations of relying solely on EXPLAIN cost estimates, preventing the system from accepting misleading or suboptimal rewrites.
+This multi-dimensional approach overcomes the limitations of relying solely on EXPLAIN cost estimates, preventing from accepting misleading or suboptimal rewrites.
 (2) Evaluation and Feedback Loop.
-Based on the generated report, the Decision agent gives a binary judge to decide whether the rewrite process should terminate or not.
+Based on the generated report, the decision agent makes a binary decision to decide whether the rewrite process should terminate or not.
 If True, the Decision agent terminates the FSM process; if False, it retrieves higher-confidence rewrite proposals from the knowledge base, then stores the current query, report, and retrieved knowledge in an agent shared buffer for the next FSM iteration.
-We deliberately avoid introducing knowledge base retrieval in the initial state, because reasoning models have input content limitations—excessive guidance information can reduce the generation freedom, potentially compromising the quality of reasoning proposal chain (Guo et al., 2025).
+We deliberately avoid introducing knowledge base retrieval in the initial state, because reasoning models have input content limitations—excessive guidance information can reduce the generation freedom, potentially compromising the quality of reasoning proposal chain (DeepSeek-r1).
 Finally, the FSM enters the Termination stage and outputs the final rewritten query along with its comprehensive optimization report.
@@
 5.1. Structured Knowledge Base
-Since DBAs encode numerous scenario-specific rewrite strategies that LLMs cannot learn from sparse examples and that fail to generalize across diverse query contexts, we consolidate these expert rules into a structured, scenario-indexed knowledge base.
-To this end, preparing the knowledge base involves the following steps: (i) collecting SQL rewrite knowledge from various resources, (ii) filtering untrustworthy content, (iii) enriching the entries, and (iv) retrieving the most relevant rules for each rewrite task.
+Many query rewrite strategies developed by DBAs cannot be effectively utilized by LLMs.
+This limitation arises because such knowledge is sparsely represented in the LLMs’ training data, making it difficult for them to extract and apply these strategies without explicit guidance provided in the prompt.
+To address this, we construct a structured knowledge base to systematically manage various rewrite strategies.
+The construction process consists of four steps: (i) collecting SQL rewrite knowledge from diverse sources, (ii) filtering unreliable content, (iii) enriching the collected entries, and (iv) retrieving the most relevant rules for each rewrite.
 The Knowledge Base $\mathcal{K}$ is defined as follows:
@@
 A Knowledge Base $\mathcal{K}$ is a set of question and answer (Q&A) pairs, denoted as $\mathcal{K}=\{\langle q_{1},a_{1}\rangle,\langle q_{2},a_{2}\rangle,\allowbreak...,\langle q_{n},a_{n}\rangle\}$.
-Each Q&A pair $\langle q_{i},a_{i}\rangle$ consists of:
-  • A question $q_{i}=\langle Text_{que_{i}},SQL_{que_{i}}\rangle$, where $Text_{que_{i}}$ describes a specific SQL rewrite issue, and $SQL_{que_{i}}$ is the original SQL.
+Each Q&A pair $\langle q_{i},a_{i}\rangle$ includes:
+  • A question $q_{i}=\langle Text_{que_{i}},SQL_{que_{i}}\rangle$, where $Text_{que_{i}}$ describes a specific SQL rewrite issue in natural language, and $SQL_{que_{i}}$ is the original SQL query that the system needs to interpret or optimize.
   • An answer $a_{i}=\langle Text_{ans_{i}},SQL_{ans_{i}}\rangle$, where $Text_{ans_{i}}$ explains the strategies used to rewrite $SQL_{que_{i}}$, and $SQL_{ans_{i}}$ is the rewritten SQL query.
@@
 Our knowledge is primarily sourced from official database documentation and the database community.
-Since rewrite strategies are often indirectly presented and scattered throughout official documents (pos, [n.d.]b; mys, [n.d.]; sql, [n.d.]; ora, [n.d.]), we employ LLMs to identify and summarize them into key points.
-We also collect knowledge from StackOverflow (sta, 2025), as it has abundant real-world query rewrite examples and insights accumulated over decades.
+Since rewrite strategies are often indirectly presented and scattered throughout official documents (postgresqldocs; mysql_docs; sqlserver_docs; oracle_db_docs), we employ LLMs to identify and summarize them into key points.
+We also collect knowledge from Stack Overflow (stackoverflow), as it has abundant real-world query rewrite examples and insights accumulated over decades.
 We retrieve all Q & A units tagged with “query rewrite” or “query optimization” that include at least one answer.
@@
 These summaries are then augmented with corroborating information from official documents.
-To do this, we embed user response summaries using Sentence Transformers (Reimers and Gurevych, 2019).
+To do this, we embed user response summaries using Sentence Transformers (sentence-transformer).
 These embeddings are then used to query a pre-indexed knowledge base of extracted points, retrieving the top three matches via cosine similarity.
@@
 Suitable matches are incorporated, enriching Q&A units with concise explanations that aid future retrieval.
+Table 2. Distribution of Knowledge UNits Across Categories
+Join Constant Predicate CTE Others
+122 2 21 20 76
 Step 4: Retrieving relevant knowledge.
-We classify these units into four categories: join optimization, constant folding, predicate simplification and others, which capture the essential aspects of query rewriting and provides a foundational structure for enabling precise retrieval of information.
-Since knowledge bases are often context-specific, we use the BM25 algorithm (Robertson et al., 2009) to retrieve the most relevant documents by computing the document relevance.
+We classify these 241 high-quality units into five categories: predicate, constant, join, CTE and others, which capture the essential aspects of query rewrite and provide a foundation for precise retrieval (Table 2).s Since knowledge is often context-specific (i.e., most rewrite strategies depend heavily on factors such as schema definitions, data distribution and available indexes), its effectiveness may diminish under changing conditions.
+In contrast, context-independent knowledge is relatively rare and involves general-purpose optimizations such as predicate simplification.
+To identify the most relevant knowledge, we use the BM25 algorithm (bm25) to retrieve the most relevant documents by computing the document relevance.
+Table 3. Constructed Hint Base
+Query Hint Grammar Description
+Hash Join /*+ HashJoin(table table[ table...]) */ Forces the use of hash joins for specified tables.
+No Nested Loop Join /*+ NoNestLoop(table table[ table...]) */ Prevents the use of nested loop joins for specified tables.
+No Merge Join /*+ NoMergeJoin(table table[ table...]) */ Prevents the use of merge joins for specified tables.
+Row Correction /*+ Rows(a b #10) */ Sets the number of rows for the join result.
+NOT_MATERIALIZE /*+ NOT_MATERIALIZE(table) */ Inlines the CTE to avoid materialization overhead.
 5.2. Hybrid SQL Corrector
@@
 Equivalence Error Correction.
-Verifying the equivalence between any pair of SQL queries is known to be an NP-hard challenge (Aho et al., 1979), and whether this problem admits a polynomial-time solution remains an open question.
-While numerous verification methods (Chu et al., 2017a; Wang et al., 2022, 2024; Ding et al., 2023) rooted in algebraic or symbolic reasoning have been proposed to prove equivalence, they typically operate on constrained query sets (e.g., specific operators, advanced SQL constructs).
-Recent studies have highlighted the potential of LLMs in deducing query equivalence with high confidence (Singh and Bedathur, 2024; Zhao et al., 2023).
+Verifying the equivalence between any pair of SQL queries is known to be an NP-hard challenge (eq_np), and whether this problem admits a polynomial-time solution remains an open question.
+While numerous verification methods (cosette; wetune; qed; sqlsolver) rooted in algebraic or symbolic reasoning have been proposed to prove equivalence, they typically operate on constrained query sets (e.g., specific operators, advanced SQL constructs).
+Recent studies have highlighted the potential of LLMs in deducing query equivalence with high confidence (singh2024exploring; LLMSQLSolverCL).
 However, relying solely on LLMs carries inherent risks.
@@
   • Tool-based Verification.
-We first use SQLSolver (Ding et al., 2023), an advanced verification tool, for an initial check.
+We first use SQLSolver (sqlsolver), an advanced verification tool, for an initial check.
 Given a pair of queries, $\langle Q_{orig},\allowbreak Q_{rewr}\rangle$, SQLSolver determines their equivalence as either equivalent, nonequivalent, or unknown (due to timeout or other limitations).
@@
 This iterative loop continues until the LLM confidently establishes query equivalence or a predefined time budget is exhausted.
-If the time budget is reached without the LLM confidently establishing equivalence, the system defaults to returning the original query ($Q_{rewri}=Q_{orig}$), ensuring that system performance is not compromised by unresolved complex rewrites.
+If the time budget is reached without the LLM confidently establishing equivalence, the system defaults to returning the original query ($Q_{rewr}=Q_{orig}$), ensuring that system performance is not compromised by unresolved complex rewrites.
 Our experiments on diverse workloads show that our method achieves higher rewrite equivalence rates than baselines (Figure 6).
 5.3. Agent Memory Buffer
-Effective context sharing is crucial in LLM-agent based query rewrite workflows.
+Effective context sharing is crucial in LLM-based workflows.
 Simply logging the entire prompt history is problematic: (1) Hallucinations due to redundant context.
-Lengthy SQL statements and redundant reasoning tokens can distract LLM, causing it to generate incorrect content (Shi et al., 2023).
+Lengthy SQL statements and redundant reasoning tokens can distract LLM, causing it to generate incorrect content (shi2023large).
 (2) Performance degradation under long-context overhead.
-Unbounded history growth incurs quadratic attention costs and increases inference latency (Jiang et al., 2023; Kuratov et al., 2024).
-To address these, we propose an agent memory buffer $\mathcal{B}$, containing memory slices =$\mathcal{B}=\{m_{1},m_{2},...,m_{n}\}$, where each slice $m_{i}$ is dedicated to a specific category of critical information (e.g., query information, rewrite proposals, retrieval knowledge).
+Unbounded history growth incurs quadratic attention costs and increases inference latency (jiang2023longllmlingua; kuratov2024babilong).
+To address these, we propose an agent memory buffer $\mathcal{B}$, containing memory slices $\mathcal{B}=\{m_{1},m_{2},...,m_{n}\}$, where each slice $m_{i}$ is dedicated to a specific category of critical information (e.g., query information, rewrite proposals).
 By extracting only essential messages across the rewrite process, our buffer curbs hallucinations and bounds context growth, thereby improving both correctness and efficiency in single-agent SQL rewrite workflows.
-Table 2. Constructed Hint Base
-Query Hint Grammar Description
-Hash Join /*+ HashJoin(table table[ table...]) */ Forces the use of hash joins for specified tables.
-No Nested Loop Join /*+ NoNestLoop(table table[ table...]) */ Prevents the use of nested loop joins for specified tables.
-No Merge Join /*+ NoMergeJoin(table table[ table...]) */ Prevents the use of merge joins for specified tables.
-Row Correction /*+ Rows(a b #10) */ Sets the number of rows for the join result.
-NOT MATERIALIZE /*+ NOT MATERIALIZE(table) */ Inlines the CTE to avoid materialization overhead.
 6. Query Hint Recommender
-Motivation.
 Query rewrite applies high-level transformations to give the DBMS optimizer a better starting point for plan generation.
 While this improved starting point is beneficial, the optimizer, which dictates the finer-grained execution operations, can still generate suboptimal plans.
-This sub-optimality arises because the optimizer’s decisions rely on cost models that are often inaccurate due to outdated statistics and simplifying assumptions (e.g., statistical independence of columns) (Stillger et al., 2001; Marcus et al., 2019; Chen et al., 2023).
-To exert more fine-grained control over plan selection, we propose embedding query hints (Marcus et al., 2021) within rewritten queries.
+This sub-optimality arises because the optimizer’s decisions rely on cost models that are often inaccurate due to outdated statistics and simplifying assumptions (e.g., statistical independence of columns) (leo; neo; base).
+To exert more fine-grained control over plan selection, we propose embedding query hints (bao) within rewritten queries.
 In this work, we introduce a dedicated hint base and use LLMs to adaptively select and incorporate effective hints into rewritten queries.
 6.1. Hint Base Construction
-Most modern database systems (e.g., PostgreSQL, MySQL, Oracle) support query hints, either inherently or by extensions (NTT OSS Center DBMS Development and Support Team, 2012; Oracle Corporation, 2024, 2019).
-In this work, we use PostgreSQL’s pg_hint_plan extension (NTT OSS Center DBMS Development and Support Team, 2012) to implement our query hint system.
-Our system supports a broader range of hints (e.g., disabling materialization and setting specific cardinality values), whereas existing works (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021) are limited to Boolean hints.
+Most modern database systems (e.g., PostgreSQL, MySQL, Oracle) support query hints, either inherently or by extensions (pg_hint_plan; mysql_hints; oracle_hints).
+In this work, we use PostgreSQL’s pg_hint_plan extension (pg_hint_plan) to implement our query hint system.
+Our system supports a broader range of hints (e.g., disabling materialization and setting specific cardinality values), whereas existing works (bao; autosteer; fastgres; steeringqo) are limited to Boolean hints.
 Selecting Existing Hints.
@@
 Introducing a Novel CTE Hint.
-To provide more precise control over Common Table Expression (CTE) execution and address associated performance issues, we introduce a new NO_MATERIALIZE hint.
-By default, PostgreSQL materializes CTEs (pos, [n.d.]a).
-This approach can lead to significant performance degradation, particularly when CTEs operate on small datasets or are executed infrequently (El-Helw et al., 2015; Zhou et al., 2007), due to the overhead of writing and reading intermediate results.
+To provide more precise control over Common Table Expression (CTE) execution and associated performance issues, we introduce a new NOT_MATERIALIZE hint.
+By default, PostgreSQL materializes CTEs (postgresqlcte), which can lead to significant performance degradation, particularly when CTEs operate on small datasets or are executed infrequently (cte1; cte2), due to the overhead of writing and reading intermediate results.
 Furthermore, the decision to materialize a CTE, as opposed to inlining it, directly impacts the optimizer’s ability to perform effective predicate pushdown into subqueries, especially after query rewrite.
-Our NO_MATERIALIZE hint directs the optimizer to inline the targeted CTE, expanding its definition into the main query as if it were a FROM-clause subquery.
+Our NOT_MATERIALIZE hint directs the optimizer to inline the targeted CTE, expanding its definition into the main query as if it were a FROM-clause subquery.
 This hint enhances query performance by eliminating materialization overhead and enabling more effective plan optimizations, such as improved predicate pushdown.
@@
 6.2. Hint Injection
-This section details our approach to using LLMs for identifying operations that could benefit from hints and then recommending appropriate ones.
-Unlike existing methods (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021) that recommend hints at the query level, where a single hint applies to all operations within that query, our strategy targets a more granular level.
-We found that query-level hints can be suboptimal, as a hint beneficial for one operation might negatively impact others.
-Therefore, we apply hints at a finer-grained level, such as to individual tables or specific operations within a query, to maximize positive impacts.
+This section explains how we use LLMs to identify operations that could benefit from hints and recommend appropriate ones.
+Prior query-level methods (bao; autosteer; fastgres; steeringqo) apply a single hint to the entire query, but we observe that such a hint can be suboptimal, as a hint that benefits one operation may degrade others.
+To overcome this limitation, our strategy instead applies hints at the operator level, such as to individual tables or specific operations within a query, to maximize their positive effects.
 Building on the validated rewrite of $Q^{\prime}$, we now turn to refining its physical execution plan through targeted hint injection.
@@
 (2) For all join operations, the LLM assesses whether the selected join operator (such as hash, merge, or nested-loop join) is suitable given the available statistics.
-Second, an LLM is then invoked to construct the most effective hints based on the obtained analysis and the query plan.
-It recommends join methods for specific tables and rectifies cardinality estimations.
-Regarding NO_MATERIALIZE, we apply this hint only if a CTE appears exactly once or processes a small, simple dataset.
+Second, the Assistant Agent processes these justifications, verifies their consistency against plan statistics, and constructs the final set of hints before forwarding them to the injection function.
+Regarding NOT_MATERIALIZE, we apply this hint only if a CTE appears exactly once or processes a small, simple dataset.
 Finally, the proposed hints are then injected into $Q^{\prime}$, producing the final optimized query $Q^{*}$.
-Our LLM-guided hint injection, unlike general existing methods, pinpoints costly operations and generates finer-grained hint recommendations.
+Unlike general existing methods (autosteer; bao; steeringqo; fastgres), our LLM-guided hint injection pinpoints costly operations and generates finer-grained hint recommendations.
 This enhances the explainability of the recommendations and leads to further improvement.
-Moreover, in contrast to ML-based hint recommendation approaches (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021) that rely on pre-trained models and multi-plan enumeration, our method requires zero training yet surpasses their performance.
+Moreover, in contrast to ML-based hint approaches (autosteer; bao; steeringqo; fastgres) that rely on pre-trained models and multi-plan enumeration, our method requires zero training yet shows a competitive performance within a small subset of query hints.
 7. Experiment
+In this section, we conduct a comprehensive evaluation of QUITE through performance comparisons, ablation studies, robustness assessments, analyses of rewrite behaviors and cost evaluations.
 7.1. Experimental Set up
-Testbed.
-Our experiments were conducted on a server featuring a 32-core Intel Xeon Platinum 8352V CPU, 251 GB of RAM, a 942 GB SSD, and PostgreSQL v14.13.
-We utilize a NVIDIA A100-SXM4 GPU with 40 GB of memory to run the rewrite process of LLM-R2.
+Testbed. Our experiments were conducted on a server featuring a 32-core Intel Xeon Platinum 8352V CPU, 251 GB of RAM, a 942 GB SSD, and PostgreSQL v14.13.
 Workloads.
-We employ three widely recognized datasets for our evaluations.
-(1) TPC-H (TPC, nd) is a well-known OLAP benchmark containing 62 columns and 22 query templates.
-We generate 63 queries from these templates, excluding Q15 due to its use of CREATE VIEW, a format unsupported by the rewrite engines of two baseline methods (Zhou et al., 2021; Li et al., 2024).
+We evaluate QUITE on standard OLAP benchmarks with long-running analytical queries, aligning with our deployment scenarios.
+We employ three widely recognized datasets and one LLM-synthesized dataset to evaluate all of the approaches.
+(1) TPC-H (TPC_H_toolkit) is a well-known OLAP benchmark containing 62 columns and 22 query templates.
+We generate 63 queries from these templates, excluding Q15 due to its use of CREATE VIEW, a format unsupported by the rewrite engines of two baseline methods (lr; llm-r2).
 We used a scale factor (SF) of 10.
-(2) DSB (Ding et al., 2021) is adapted from the TPC-DS benchmark (Nambiar and Poess, 2006) and is notable for its complex data distribution and challenging long-context query templates, which consists of 52 query templates.
+(2) DSB (ding2021dsb) is adapted from the TPC-DS benchmark (tpcds) and is notable for its complex data distribution and challenging long-context query templates, which consists of 52 query templates.
 We generate 156 queries based on these templates with $SF=10$.
-(3) Calcite (apa, 2021) is a real-world workload used to evaluate rewrite rules.
-We randomly select 58 queries and generate 10G of data with uniform distributions following previous work (Dong et al., 2023).
+(3) Calcite (apache_calcite_test) is a real-world workload used to evaluate rewrite rules in Apache Calcite (apache).
+We randomly select 58 queries and generate 10G of data with uniform distributions following previous work (slabcity).
+(4) StackOverflow (StackOverflow-Math) (stackoverflow-math) is a 13.8 GB dataset from the Mathematics community on the Stack Exchange network [^1] Stack Overflow is the flagship site of the network (sqlstorm), containing math-related posts and metadata as of October 3, 2024.
+Following  (sqlstorm), we generate queries using predefined complexity prompts (P1–P5) and obtain 43 valid queries after dialect correction.
 Baselines. We compare our system against the following state-of-the-art query rewrite baselines:
-  • LearnedRewrite (LR) (SIGMOD 2022)  (Zhou et al., 2021) uses a Monte Carlo Tree Search (MCTS) algorithm with a learned cost estimation model to explore a SQL policy tree of rewrite rule orders.
+  • LearnedRewrite (LR) (SIGMOD 2022)  (lr) uses a Monte Carlo Tree Search (MCTS) algorithm with a learned cost estimation model to explore a SQL policy tree of rewrite rule orders.
 We adopt the original implementation and use the cost estimation model provided in their official code repository.
-  • LLM-R2 (VLDB 2025) (Li et al., 2024) is a rule-based query rewrite system built on Apache Calcite (Begoli et al., 2018).
+  • LLM-R2 (VLDB 2025) (llm-r2) is a rule-based query rewrite system built on Apache Calcite (apache).
 It uses LLM’s In-Context Learning (ICL) ability to select query rewrite rules according to high quality demonstration queries.
-  • R-bot (Sun et al., 2024) is an LLM-based system that leverages an embedded retrieval-augmented generation (RAG) knowledge base and reflective reasoning to select promising rewrite rules.
+  • R-bot (VLDB 2025) (r-bot) is an LLM-based system that leverages an embedded retrieval-augmented generation (RAG) knowledge base and reflective reasoning to select promising rewrite rules.
   • LLM Agent rewrites queries using an individual LLM agent. To ensure a fair comparison, this LLM agent can access target database statistics and our SQL corrector for self-reflection.
-Table 3. Different Methods on Query Latency (s)
-Methods TPC-H (SF=10) DSB (SF=10) Calcite (SF=10)
-Mean Median 75th 95th Mean Median 75th 95th Mean Median 75th 95th
-Original 69.84 9.64 32.75 300.00 32.62 4.85 10.29 300.00 23.88 2.58 8.64 122.36
-LearnedRewrite 37.57 9.97 30.55 202.31 31.93 5.14 15.18 251.42 22.88 2.32 8.09 122.58
-LLM-R2(GPT-4o) 57.09 9.30 30.50 300.00 10.53 4.11 8.36 32.88 16.98 2.11 7.75 50.97
-LLM-R2(Claude-3.7) 57.68 9.89 30.07 300.00 9.11 3.82 8.24 27.48 16.96 2.55 8.19 50.07
-LLM-R2(DS-R1) 56.73 9.25 29.33 300.00 9.16 3.84 8.21 24.21 16.90 2.05 7.71 49.91
-R-Bot (GPT-4o) 32.64 9.48 26.25 172.20 21.75 4.26 8.80 119.19 22.21 2.34 8.35 122.14
-R-Bot (Claude-3.7) 33.89 9.71 29.32 195.51 19.89 4.27 8.99 112.24 23.38 2.60 8.64 122.95
-R-Bot (DS-R1) 33.70 9.88 28.38 176.90 21.21 4.44 9.11 113.91 22.71 2.52 8.43 122.89
-LLM Agent (GPT-4o) 69.03 10.00 27.02 300.00 33.44 4.39 9.65 300.00 15.47 2.03 7.32 35.34
-LLM Agent (DS-R1) 42.79 9.94 27.95 300.00 13.53 3.87 7.93 30.18 15.18 1.65 7.55 36.70
-LLM Agent (DS-V3) 60.12 9.22 29.92 300.00 32.34 4.19 9.99 300.00 16.49 2.53 7.72 45.15
-LLM Agent (Claude-3.7) 57.46 11.04 36.99 300.00 25.32 4.38 8.84 300.00 15.67 1.59 7.68 44.54
+Table 4. Query Latency of Different Methods (s)
+Methods TPC-H (SF=10) DSB (SF=10) Calcite (SF=10) StackOverflow (SF=10)
+Mean Median 75th 95th Mean Median 75th 95th Mean Median 75th 95th Mean Median 75th 95th
+Original 69.84 9.64 32.75 300.00 32.62 4.85 10.29 300.00 23.88 2.58 8.64 122.36 48.02 9.47 20.80 300.00
+LearnedRewrite 37.57 9.97 30.55 202.31 31.93 5.14 15.18 251.42 22.88 2.32 8.09 122.58 46.25 9.43 18.35 300.00
+LLM-R2(GPT-4o) 57.09 9.30 30.50 300.00 10.53 4.11 8.36 32.88 16.98 2.11 7.75 50.97 47.08 9.19 16.06 300.00
+LLM-R2(Claude-3.7) 57.68 9.89 30.07 300.00 9.11 3.82 8.24 27.48 17.80 2.30 7.83 57.28 46.97 8.40 16.30 300.00
+LLM-R2(DS-R1) 56.73 9.25 29.33 300.00 9.16 3.84 8.21 24.21 16.90 2.05 7.71 49.91 43.42 9.21 14.14 300.00
+R-Bot (GPT-4o) 32.64 9.48 26.25 172.20 21.75 4.26 8.80 119.19 22.21 2.34 8.35 122.14 46.43 9.37 16.02 300.00
+R-Bot (Claude-3.7) 33.89 9.71 29.32 195.51 19.89 4.27 8.99 112.24 23.38 2.60 8.64 122.95 52.06 9.52 19.25 300.00
+R-Bot (DS-R1) 33.70 9.88 28.38 176.90 21.21 4.44 9.11 113.91 22.71 2.52 8.43 122.89 50.57 9.46 17.39 300.00
+LLM Agent (GPT-4o) 69.03 10.00 27.02 300.00 33.44 4.39 9.65 300.00 15.47 2.03 7.32 35.34 50.90 7.96 26.73 300.00
+LLM Agent (DS-R1) 42.79 9.94 27.95 300.00 13.53 3.87 7.93 30.18 15.18 1.65 7.55 36.70 27.14 6.38 9.40 272.55
+LLM Agent (DS-V3) 60.12 9.22 29.92 300.00 32.34 4.19 9.99 300.00 16.49 2.53 7.72 45.15 54.92 10.33 23.60 300.00
+LLM Agent (Claude-3.7) 57.46 11.04 36.99 300.00 25.32 4.38 8.84 300.00 15.67 1.59 7.68 44.54 36.23 7.58 21.86 283.42
 […]
@@
 […]
-(a) TPC-H Benchmark
-(b) DSB Benchmark
-(c) Calcite Benchmark
+(a) TPC-H
+(b) DSB
+(c) Calcite
+(d) StackOverflow
 Figure 6. Rewrite Equivalence and Improvement Numbers on Different Benchmarks
+Figure 7. Comparison of Overall Latency (h)
+Comparison with GenRewrite (genrewrite).
+GenRewrite explores query rewriting with natural-language rewrite rules and iterative correction, but the lack of released code prevents direct comparison.
+At a high level, it accumulates natural-language rewrite knowledge to guide subsequent LLM-based rewriting.
+In contrast, QUITE formulates query rewrite as a decision process and directly optimizes execution through iterative refinement with database feedback, enabling more robust and data-aware optimizations.
 QUITE Settings.
-We use DeepSeek-R1 as the reasoning agent due to its strong analysis capability, and Claude-3.7-Sonnet as the rewrite, decision and assistant agents as default.
-A temperature of 0 is applied to all LLMs to ensure output stability  (DeepSeek-AI, 2025; Anthropic, 2025).
+We use DeepSeek-R1 as the reasoning agent due to its strong analysis capability, and Claude-3.7-Sonnet as the rewrite, decision, and assistant agents by default.
+A temperature of 0 is applied to all LLMs to ensure output stability (DeepSeek_doc; anthropic_doc).
 The number of iterations is limited to 2 to prevent excessive FSM rewrite cycles.
@@
 Query Evaluation Approach.
-Before each query, we restart the evaluation database to clear all caches.
-The query execution process involves an initial warm-up run (to mitigate cold start impacts), followed by three measured executions.
+Before each query, the database is restarted to clear caches.
+The query execution process involves an initial warm-up run to mitigate cold start impacts, followed by three measured executions.
 We report the average execution time of these three runs.
-Queries that run for more than 300 seconds are terminated and their time is recorded as 300 seconds (Li et al., 2024), to reflect impractical execution durations.
-If a rewritten query produces different output than the original or contains syntax errors, its equivalence metric is marked as False.
-To simulate a realistic query optimization scenario when dealing with such nonequivalent queries, we assign the rewritten query the execution time of its original counterpart.
+Queries exceeding 300 seconds are terminated and recorded (llm-r2; r-bot), with capped values included in all latency statistics (mean, median, 75th, and 95th percentiles).
+Rewritten queries with syntax errors or output mismatches are marked as non-equivalent.
+In such cases, the execution time of the original query is used.
+If a rewrite times out, we recheck equivalence on a smaller scale: non-equivalent rewrites fall back to the original, while equivalent rewrites are executed with capped latency.
 7.2. Performance Comparison
 Query Execution Latency.
-As presented in Table 3, queries optimized by QUITE demonstrate significantly reduced execution times, consistently outperforming all baseline methods.
-Across the TPC-H, DSB, and Calcite benchmarks, QUITE achieves substantial average execution time reductions.
-Specifically, QUITE reduces execution time by 31.9%, 81.7%, and 56.4% compared to LR; 54.9%, 35.8%, and 41.0% over LLM-R2; and 40.2%, 56.8%, and 34.3% against LLM Agent.
-When evaluated against R-Bot, QUITE achieves reductions of 21.6%, 70.6%, and 55.1%.
-QUITE’s advantages are further illustrated by the rewritten improvement rates shown in Figure 6, where QUITE achieves the highest ratios of 31.7%, 37.8%, and 34.5% on TPC-H, DSB, and Calcite, respectively, indicating strong generalization capabilities across different workloads.
-The optimization benefits of QUITE over baselines are more pronounced on complex datasets.
-Notably, the improvement margin of QUITE over LR is substantially larger for the DSB dataset compared to the TPC-H and Calcite datasets.
-This is primarily because: (1) The LR we use is trained on the TPC-H, LR can proficiently identify effective rewrite rules and navigate their constrained search spaces.
-This existing proficiency limits the remaining optimization potential, thereby reducing the margin for QUITE to demonstrate substantial further improvements.
-(2) LR employs a greedy search within its MCTS algorithm, which proves less effective in the expansive search spaces of complex datasets like DSB, unlike its performance on simpler datasets (e.g., Calcite) with smaller search spaces and inherently less optimization potential.
+As shown in Table 4, queries optimized by QUITE demonstrate significantly reduced execution times, consistently outperforming all baseline methods.
+Across the TPC-H, DSB, Calcite and StackOverflow benchmarks, QUITE achieves substantial average execution time reductions.
+Specifically, QUITE reduces execution time by 31.9%, 81.7%, 56.4% and 72.7% compared to LR; 54.9%, 35.8%, 41.0% and 71.0% over LLM-R2; and 40.2%, 56.8%, 34.3% and 53.5% against LLM Agent.
+When evaluated against R-Bot, QUITE achieves reductions of 21.6%, 70.6%, 55.1% and and 72.8%.
+QUITE’s advantages are further illustrated by the rewritten improvement rates shown in Figure 6, where QUITE achieves the highest ratios of 31.7%, 37.8%, 34.5% and 65% on TPC-H, DSB, and Calcite, respectively, indicating strong generalization capabilities across different workloads.
+The optimization benefits of QUITE are more pronounced on complex datasets.
+Notably, on the LLM-generated StackOverflow workload, we did not curate queries to disadvantage baselines.
+Since many LLM-generated queries use broad SQL dialects that are not directly supported by the strict syntax of Apache Calcite used in rule-based pipelines (lr; llm-r2; r-bot), we minimally normalized only such dialect mismatches (e.g., LISTAGG vs.
+STRING_AGG), discarding unresolvable cases.
+Under this controlled setup, rule-based pipelines remain bottlenecked by dialect constraints and limited rule and pattern coverage on complex structures, while QUITE is more robust and delivers larger latency gains.
+Table 5. Unified Ablation Study Results on DSB (SF=10)
+Metrics Structured Knowledge Base Different LLMs Hint Construction Methods
+w/o Q&A Units Raw Q&A Units Filtered Q&A Units DS-R1 + GPT-4o DS-R1 + DS-V3 DS-R1 + Claude-3.7 DS-V3 + Claude-3.7 Claude-3.7 + Claude-3.7 w/o Hints Bao’s Hint Base Full Hints (no GUC) Ours + Scan Our Hint Base
+Mean 9.96 9.10 6.08 31.38 28.07 6.08 23.41 12.76 6.08 6.93 7.06 6.04 5.85
+Median 3.94 3.85 3.78 4.64 4.43 3.78 4.32 4.01 3.78 4.65 4.02 3.89 3.35
+75th 8.49 8.10 7.98 9.56 9.07 7.98 9.64 8.74 7.98 8.46 8.64 8.06 7.63
+95th 24.81 23.01 21.43 300.00 300.00 21.43 148.47 58.38 21.43 20.41 22.52 20.29 19.97
 Figure 8. Ablation Study of Hint Injection on DSB Benchmark
-Rewritten Equivalence.
-Figure 6 details the number of equivalent and improved query rewrites produced by each method across the datasets.
-Notably, direct rewrites from the LLM Agent exhibit significant semantic equivalence issues.
-In contrast, QUITE achieves substantially higher equivalence rates–attaining 100%, 96.8%, and 98.3% on TPC-H, DSB, and Calcite, respectively.
-This superior performance is attributable to the multi-stage correction mechanisms integrated into QUITE’s iterative FSM-based process.
 Figure 9. Ablation Study of Hybrid SQL Corrector
-7.3. Ablation Study
-We design our ablation study from four perspectives by selectively replacing or removing system components.
-Rewrite Middleware.
-To assess the hybrid SQL corrector, we removed it and evaluated its impact on query equivalence and execution efficiency using the DSB dataset.
-As shown in Figure 8, judge accuracy measures the proportion of correct executions among all decisions made by the tool or LLM while coverage ratio indicates the proportion of queries for which the tool or LLM was able to make a definitive decision.Its removal significantly degrades performance: nonequivalent queries increase from 5 to 11, unchanged queries by 13, and improved queries drop by 17.
-These results highlight the corrector’s essential role in ensuring query quality and efficiency.
-We also evaluate the impact of the structured knowledge component.
-As shown in Table 4, QUITE consistently outperforms its variant without this module across all metrics.
-The improvement stems from high-quality Q&A pairs and domain knowledge embedded in the knowledge base, which effectively guide more optimized rewrites.
-Table 4. Ablation Study of Structured Knowledge Base
-Methods DSB (SF=10)
-Mean Median 75th 95th
-Without Knowledge 9.96 3.94 8.49 24.81
-With Knowledge 6.08 3.78 7.98 21.43
 Table 6. Ablation Study of the Reasoning Agent and FSM
@@
 6.08 3.78 7.98 21.43
+Overall Latency.
+As shown in Figure 7, we evaluate overall execution time on TPC-H (SF=50) and DSB (SF=50) with a per-query time limit of 8000 seconds, using Claude-3.7-Sonnet across all baselines.
+On TPC-H and DSB, QUITE reduced execution time by 71.4% and 53.6% compared with the original queries, outperforming LR (47.6%, 53.4%), LLM-R2 (71.9%, 20.0%), and R-Bot (29.1%, 45.5%), respectively.
+These results demonstrate QUITE’s strong end-to-end advantage.
+QUITE’s rewrite time on the DSB dataset is relatively longer, which can be attributed to two main factors: first, DeepSeek-R1 requires more time to process difficult problems; second, for complex queries, the decision agent often performs an additional iteration to explore further optimizations.
+Rewritten Equivalence.
+Figure 6 details the number of equivalent and improved rewrites produced by each method across the datasets.
+Notably, direct LLM-agent rewrites, especially by Claude-3.7-Sonnet on StackOverflow, often violate equivalence due to subtle semantic drift during query restructuring.
+In contrast, QUITE achieves substantially higher equivalence rates, attaining 100%, 96.8%, 98.3% and 90.7% on TPC-H, DSB, Calcite and StackOverflow.
+In addition, rule-based methods suffer from incomplete semantic equivalence, as many rules in their rewrite engines (e.g., Apache Calcite (apache)) lack formal verification due to the inherent difficulty of equivalence checking (wetune; sqlsolver).
+Despite this, our approach achieves the highest equivalence rate across all benchmarks, advancing the state of the art.
+7.3. Ablation Study
+We design our ablation study from four perspectives by selectively replacing or removing system components.
+Rewrite Middleware.
+First, we ablated the hybrid SQL corrector to measure its effect on equivalence and execution.
+Figure 9 reports Judge Accuracy (the fraction of correct decisions) and Coverage Ratio (the fraction of queries with a definitive decision).
+Its removal significantly degrades performance: nonequivalent queries increase from 5 to 11, unchanged queries increase by 13, and improved queries drop by 17.
+These results highlight the corrector’s essential role in ensuring query quality and efficiency.
+We also evaluate the impact of the structured knowledge base.
+As shown in Table 5, QUITE consistently outperforms its variant without this module across all metrics.
+The improvement stems from high-quality domain knowledge embedded in the knowledge base, which effectively guide more optimized rewrites.
+To justify the LLM-based filtering, we compare the 241 filtered units against all 3,432 collected units.
+The filtered version delivers significantly stronger execution performance: five queries achieve a $2\times$$5\times$ speedup, and more than eight queries exceed $5\times$.
+This confirms that consensus-aligned, high-quality knowledge provides significantly more effective guidance for LLM rewriting, enabling stable and efficient query optimization.
+Table 7. Robustness Study on TPC-H and DSB Benchmarks under Different Scale Factors (SF)
+Methods TPC-H DSB
+SF=1 SF=10 SF=30 SF=1 SF=10 SF=30
+Mean Med 75th 95th Mean Med 75th 95th Mean Med 75th 95th Mean Med 75th 95th Mean Med 75th 95th Mean Med 75th 95th
+Original 29.48 0.69 1.20 300.00 69.31 9.42 30.08 300.00 64.76 20.61 46.69 300.00 7.22 0.27 0.69 13.19 32.62 4.85 10.29 300.00 39.42 5.06 18.22 300.00
+LearnedRewrite 1.46 0.52 1.10 3.91 37.57 9.97 30.55 202.31 52.59 18.90 43.50 300.00 17.15 0.42 1.52 81.14 31.93 5.14 15.18 251.42 43.88 7.09 20.73 300.00
+LLM-R2(DS-R1) 15.38 0.69 1.22 7.69 56.73 9.25 29.33 300.00 54.67 18.30 58.11 300.00 4.30 0.23 0.57 2.15 9.16 3.84 8.21 24.21 23.01 5.01 11.87 87.15
+R-Bot (GPT-4o) 1.49 0.66 1.03 6.38 32.64 9.48 26.25 172.20 54.54 19.85 61.12 300.00 9.06 0.25 0.63 21.89 21.75 4.26 8.80 119.19 31.93 5.15 16.27 300.00
+QUITE (Rw. on SF=1) 1.02 0.64 1.01 3.79 36.87 13.53 29.02 279.51 54.36 25.02 43.03 289.48 2.25 0.24 0.50 2.74 10.58 4.18 8.59 30.79 23.23 4.67 12.91 111.83
+QUITE (Rw. on SF=10) 1.08 0.68 1.22 3.79 26.06 9.28 24.19 75.64 45.90 17.44 57.26 155.72 2.82 0.24 0.51 1.72 6.08 3.78 7.98 21.43 17.36 4.34 9.28 77.62
+QUITE (Rw. on SF=30) 5.72 0.50 1.17 5.17 30.50 9.96 24.29 96.68 40.65 16.89 37.96 131.97 6.65 0.26 0.56 3.20 8.74 3.94 8.01 23.15 16.35 4.45 9.26 74.33
 LLM-Agent-based Query Rewrite FSM.
-To evaluate the impact of the FSM framework within our system, we performed two modifications: (1) the loop responsible for equivalence checking was removed, and (2) the decision stage was eliminated.
-These changes transformed the process into a linear, end-to-end pipeline while retaining all components and agent tools involved in query rewriting.
-At the same time, to further assess the impact of the MDP-based decision-making on the Reasoning agent’s capabilities, we conducted an ablation study by removing the MDP component.
-As shown in Table 5, QUITE outperforms its variant across all evaluation metrics.
-Notably, the average execution time with the MDP is reduced by 52.2% compared to the version without it, demonstrating that the MDP significantly enhances the effectiveness of the Reasoning agent.
-Table 6. Ablation Study of QUITE using Different LLMs
-Methods DSB (SF=10)
-Mean Median 75th 95th
-Deepseek-R1 + GPT-4o 31.38 4.64 9.56 300.00
-Deepseek-R1 + Deepseek-V3 28.07 4.43 9.07 300.00
-Deepseek-R1 + Claude-3.7 6.08 3.78 7.98 21.43
-Table 7. Ablation Study of Hint Construction Methods
-Methods DSB (SF=10)
-Mean Median 75th 95th
-QUITE (without hints) 6.08 3.78 7.98 21.43
-QUITE (with Bao’s Hint Base) 6.93 4.65 8.46 20.41
-QUITE (with Our Hint Base) 5.85 3.35 7.63 19.97
-Table 8. Robustness Study of Different Data Sizes
-Methods TPC-H (SF=1) TPC-H (SF=10) TPC-H (SF=30)
-Mean Median 75th 95th Mean Median 75th 95th Mean Median 75th 95th
-Original 29.48 0.69 1.20 300.00 69.31 9.42 30.08 300.00 64.76 20.61 46.69 300.00
-LearnedRewrite 1.46 0.52 1.10 3.91 37.57 9.97 30.55 202.31 52.59 18.90 43.50 300.00
-LLM-R2(Deepseek-R1) 15.38 0.69 1.22 7.69 56.73 9.25 29.33 300.00 54.67 18.30 58.11 300.00
-R-Bot (GPT-4o) 1.49 0.66 1.03 6.38 32.64 9.48 26.25 172.20 54.54 19.85 61.12 300.00
-QUITE 1.02 0.64 1.01 3.79 26.06 9.28 24.19 75.64 40.65 16.89 37.96 131.97
-Table 9. Data-Aware Analysis on TPC-H Benchmark
-Methods TPC-H (SF=1) TPC-H (SF=10) TPC-H (SF=30)
-Mean Median 75th 95th Mean Median 75th 95th Mean Median 75th 95th
-Original 29.48 0.69 1.20 300.00 69.31 9.42 30.08 300.00 64.76 20.61 46.69 300.00
-QUITE (Rewrite on SF=1) 1.02 0.64 1.01 3.79 36.87 13.53 29.02 279.51 54.36 25.02 43.03 289.48
-QUITE (Rewrite on SF=10) 1.08 0.68 1.22 3.79 26.06 9.28 24.19 75.64 45.90 17.44 57.26 155.72
-QUITE (Rewrite on SF=30) 5.72 0.50 1.17 5.17 30.50 9.96 24.29 96.68 40.65 16.89 37.96 131.97
+To evaluate the impact of the FSM framework, we make two modifications: (1) remove the equivalence-checking loop and (2) remove the decision stage.
+This converts the process into a linear pipeline while keeping all components and agent tools for rewriting.
+To further isolate the contribution of the MDP-based reasoning agent, we also conduct an ablation that removes the MDP component.
+As shown in Table 6, QUITE outperforms all variants across metrics.
+Notably, adding the MDP reduces the average execution time by 52.2% compared to the version without it, indicating that the MDP substantially improves the reasoning agent’s effectiveness.
 Query Hint Injection.
 We evaluate the impact of hint injection on queries rewritten by QUITE.
-As shown in Table 3, the insertion of hints leads to a further decrease in average execution times by 1.8%, 3.8%, and 0.3% on TPC-H, DSB, and Calcite, respectively.
-These performance gains demonstrate that appropriate hints enable additional fine-grained optimization, even when queries have already been extensively optimized through rewriting.
-To highlight the advantage of our LLM-based hint injection approach, we compare its performance on rewritten queries against using LLMs to recommend Bao’s candidate hint sets.
-As shown in Table 7 and Figure 7, within the hint space defined at the GUC level provided by Bao, the overall trend indicates degraded performance.
-Furthermore, Figure 7 presents a comparison of execution times for 24 positively optimized queries on DSB, both before and after hint injection, as well as against the original queries.
-Notably, some queries that exhibited degraded performance after rewriting were able to exceed their original execution efficiency through proper hint injection.
-This demonstrates that hint injection can effectively recover and enhance performance in cases where the initial rewrites were suboptimal.
-Base Model Selection.To evaluate the role of the base models within our framework, we conducted experiments using different model combinations.
-DeepSeek-R1 was retained as the Reasoning agent due to its low cost and strong chain-of-thought reasoning capabilities, making it indispensable for this role.
-For the Decision agent, we tested three alternatives: Claude-3.7-Sonnet, DeepSeek-V3, and GPT-4o.
-As shown in Table 6, rewrites using Claude-3.7-Sonnet as the Decision agent significantly outperform those from the other two models.
-This is due to its strengths in summarization and decision-making, which require long-context understanding, precise reasoning, and accuracy—areas where Claude-3.7-Sonnet excels.
-In contrast, DeepSeek-V3 shows higher hallucination rates, and GPT-4o offers lower accuracy and weaker long-context handling, making both less suitable for our framework.
+As shown in Table 4, hint injection further reduces average execution time by 1.8%, 3.8%, and 0.3% on TPC-H, DSB, and Calcite, respectively, indicating that appropriate hints enable additional fine-grained optimization even after extensive rewriting.
+To highlight the advantage of our approach, we compare it with using LLMs to select Bao’s GUC-level hint sets.
+As shown in Table 5 and Figure 8, this baseline exhibits an overall performance degradation within Bao’s predefined hint space.
+Furthermore, Figure 8 shows execution times for 24 positively optimized DSB queries before and after hint injection, as well as for the original queries.
+In several cases, hint injection recovers and surpasses the original performance when rewriting alone is suboptimal.
+This demonstrates hint injection can recover and enhance performance in cases where the initial rewrites were suboptimal.
+Figure 10. Rewrite Categories for Different Benchmarks
 Table 8. Cost Analysis on TPC-H Benchmark
 Methods Average Time (s) Cost (\$)
-LLM-R2 (Claude-3.7) 399.30 0.93
-R-Bot (Claude-3.7) 256.96 20.94
-QUITE (DS-R1 + Claude-3.7) 291.59 6.76
+LLM-R2 (Claude-3.7) 391.23 0.016
+R-Bot (Claude-3.7) 118.33 0.417
+QUITE (DS-R1 + Claude-3.7) 218.25 0.200
+Sensitivity Study.
+We examine QUITE’s sensitivity to the number of rewrite rounds, hint thresholds, and different LLM models.
+We first test different model combinations.
+We fix DeepSeek-R1 as the reasoning model and vary the other models among Claude-3.7-Sonnet, DeepSeek-V3, and GPT-4o, and then we replace the reasoning model with DeepSeek-V3 and Claude-3.7-Sonnet.
+As shown in Table 5, the combination of DeepSeek-R1’s strong reasoning capabilities and Claude-3.7-Sonnet’s long-context understanding significantly outperforms all other combinations.
+For hint thresholds, as shown in Table 5, we test the effect of adding the scan method to our hint base and the effect of using all hints (excluding GUC).
+The former yields an execution time nearly identical to that of the rewritten query, while the latter causes a slight performance degradation.
+This suggests that the size of our current hint search space is appropriate.
 7.4. Robustness Study
-We assess the robustness of QUITE by testing stability across data scales.
-We scale the TPC-H dataset from SF = 10 to SF = 1 and SF = 30, then execute rewritten queries on all three versions.
-Table 8 shows that QUITE delivers significant optimizations at each scale and outperforms all baselines.
-In contrast, LearnedRewrite and LLM-R2 often produce rewrites that worsen performance for complex queries.
-We evaluate the robustness of the QUITE framework by analyzing its stability.
-The stability assessment demonstrates that QUITE generalizes effectively across diverse data scales.
-To evaluate the stability of QUITE across varying database sizes, we extended the primary TPC-H $SF=10$ dataset by generating proportionally scaled versions at scale factors of $SF=1$ and $SF=30$.
-For each evaluated method, queries rewritten based on the $SF=10$ dataset were executed and timed on the $SF=1$, $SF=10$, and $SF=30$ datasets.
-As shown in Table 8, QUITE consistently achieved substantial optimization improvements across all scales, outperforming all baseline methods.
+Scale-Aware Robustness.
+We evaluate QUITE on TPC-H and DSB under scale factors $SF\in\{1,10,30\}$, where rewriting and execution are performed on the same scale.
+As Table 7 shows, QUITE consistently achieves substantial latency reductions at every scale.
+In contrast, LR, LLM-R2, and R‑Bot often degrade as the data scale changes.
+These results demonstrate that QUITE maintains stable efficiency under varying data volumes and adapts robustly to distribution shifts within each benchmark.
+Cross-Scale Robustness.
+To examine generalization across data distributions, we perform cross-scale evaluation: rewrites produced at one scale ( $SF\in\{1,10,30\}$) are executed on all others.
+As Table 7 shows, QUITE achieves its strongest gains when rewrite and execution scales match, while mismatched cases show only minor regressions and still outperform all baselines.
+This behavior highlights QUITE’s strong data adaptivity and robustness, effectively learning to tailor rewrites to underlying data characteristics without overfitting to a specific workload or scale.
 7.5. Further Analysis
-Data-Aware Analysis.
-We evaluated the sensitivity of QUITE using a cross-testing strategy.
-Specifically, QUITE generated three sets of rewritten queries, each optimized for the data statistics of a specific TPC-H scale: $SF$=1, $SF$=10, and $SF$=30.
-Each set was subsequently executed across all three datasets.
-As shown in Table 9, the best performance consistently occurred when queries were executed on the dataset for which they were originally optimized.
-This result confirms that QUITE effectively distinguishes between different data distributions and produces rewrites that are well-adapted to the target datasets.
+Rewrite Type Analysis.
+We report successful rewrites that achieve over a 10% performance improvement while maintaining equivalence across four benchmarks.
+As shown in Figure 10, most gains come from join, CTE, and predicate rewrites, while the rest are benchmark-specific.
+Compared with rule-based baselines (lr; llm-r2; r-bot), QUITE produces deeper rewrites beyond pattern matching.
+We further analyse efficient rewrite types, query structures, Q&A units, and query hints in our Github files (quite_ap; quite_rt).
 Cost Analysis.
-As shown in Table 10, LLM-R2 incurs high time expenditure due to demonstration construction, while R-Bot consumes substantial tokens during stepwise rewriting.
+We conducted a cost analysis experiment on 63 queries of TPC-H.
+As shown in Table 8, LLM-R2 incurs high time cost for demonstration construction, while R-Bot consumes substantial tokens for stepwise rewriting.
 In contrast, QUITE offers a balanced trade-off between time expenditure and cost.
-By generating much of its output with DeepSeek-R1, it greatly reduces monetary cost.
+By generating much of its output with DeepSeek-R1, QUITE greatly reduces monetary cost.
 Additionally, it avoids pre-training and extensive knowledge base construction.
-To further control execution time, we cap FSM iterations, ensuring reliable output quality without incurring excessive time expenditure.
 8. Conclusion
-This paper presents QUITE, a training-free and feedback-aware system that smartly leverages LLMs to rewrite SQL queries into semantically equivalent forms with significantly improved performance.
-We rethink the query rewrite process and identify unique opportunities to support a wider range of query patterns and rewrite strategies with good scalability.
-The implemented QUITE  system significantly outperforms state-of-the-art methods in both query performance and query coverage on widely used benchmarks.
+This paper presents QUITE, a training-free and feedback-aware system that smartly leverages LLM Agents to rewrite SQL queries into semantically equivalent forms with significantly improved performance.
+We rethink the query rewrite process and identify unique opportunities to support a wider range of query patterns and rewrite strategies with strong scalability.
+The implemented QUITE  system significantly outperforms state-of-the-art methods in both query performance and query coverage on widely used benchmarks and synthetic workloads.